### v4

# Computing Generalized Ranks of Persistence Modules via Unfolding to Zigzag Modules

Abstract
For a persistence module $\mathbb{M}$ indexed by a finite poset $P$, its generalized rank is defined as the rank of the limit-to-colimit map for the diagram of constituent vector spaces and linear maps of $\mathbb{M}$ over the poset $P$. For $2$-parameter persistence modules, a zigzag persistence based algorithm has been proposed that takes advantage of the fact that generalized rank for $2$-parameter modules is equal to the number of full intervals in a zigzag module defined on the ‘boundary’ of the poset. Analogous definition of boundary for $d$-parameter persistence modules or general poset indexed persistence modules does not seem plausible. To overcome this difficulty, we first unfold a given module $\mathbb{M}$ into a zigzag module $\mathbb{M}_{ZZ}$ and then check how many full interval modules in a decomposition of $\mathbb{M}_{ZZ}$ can be folded back to remain full in a decomposition of $\mathbb{M}$. This number determines the generalized rank of $\mathbb{M}$. We give an efficient algorithm to compute this number for the case when $\mathbb{M}$ is induced by a filtration of a simplicial complex over a finite poset. For special cases where $\mathbb{M}$ is the degree-$d$ homology functor on a filtration of a $d$-complex, we obtain a more efficient algorithm including a linear time algorithm for degree-$1$ homology for graphs.
1 Introduction
It is well known that one-parameter persistence modules decompose into interval modules whose supports constitute persistence diagram, or equivalently barcode of the given module, a fundamental object of study in topological data analysis [6, 17, 19, 33]. There are many situations in which persistence modules are parameterized over more than one parameter [5, 9, 20, 27, 31]. Unfortunately, such multiparameter persistence modules do not necessarily admit a nice decomposition into intervals only. Instead, they may decompose into indecomposables that are more complicated [9]. To overcome this difficulty, inspired by the work of Patel [34], Kim and Mémoli [25] proposed a decomposition of poset-indexed modules (satisfying some mild condition) into signed interval modules. In analogy to the one-parameter case, the supports of these signed interval modules with multiplicity are called the signed barcode of the given module [3]. The multiplicity of the intervals is given by the Möbius inversion of a rank-invariant function; see [25, 34]. Botnan, Oppermann, and Oudot [3] recently showed that a unique minimal signed barcode of a given persistence module in terms of rectangles can be computed efficiently and raised the question of efficient computation of other types of signed barcodes or invariants; see the recent work in [12] for more exposition along this direction.
At the core of computing these signed barcodes for a persistence module $\mathbb{M}$ is the problem of computing the generalized rank for an interval $I$ that is defined as the rank of the limit-to-colimit map for the diagram of vector spaces of the restricted module $\mathbb{M}|_{I}$; see [25]. Recently, Dey, Kim, Mémoli [16] showed that, for a $2$-parameter persistence module, this generalized rank is given by the number of full intervals in the decomposition of a zigzag module $\mathbb{M}_{ZZ}$ that is a restriction of $\mathbb{M}$. This becomes immediately useful because there are efficient algorithms known for computing the barcode of a zigzag module  [14, 15, 30, 32]. However, this result is limited to $2$-parameter persistence modules because the zigzag module needs to be defined on the boundary of a “two dimensional” interval. Beyond the $2$-parameter, this boundary does not remain to be a path and hence poses a challenge in defining an appropriate zigzag module.
In this paper, we address the above problem and present an algorithm to compute generalized rank efficiently for finite-dimensional modules indexed by finite posets. The approach uses the idea of straightening up the input persistence module into a module defined over a zigzag path. We call the process unfolding the module. We compute a decomposition of the resulting zigzag module into interval modules using a known algorithm. Then, we design an algorithm that aims to fold the full interval modules (supported on entire zigzag path) in the decomposition of the zigzag module back to the original module. The ones which fold successfully to a full interval summand (supported on the entire poset) in the original module gives us the generalized rank according to a result of Chambers and Letscher [11].
A viable approach to compute the generalized rank would be to compute the limit and colimit of a given persistence module separately, say with the recent algorithm in [35], and then compute the rank of the limit-to-colimit map. Finding an efficient implementation of this approach remains open. Each of the computations for limit, colimit, and then the rank of the map between them may incur considerable computational cost in practice. Our approach also has three distinct computations, unfolding the module, then computing a zigzag persistence followed by a folding process. Among these, the first one is done by a simple graph traversal, the second one can be done with recent efficient practical zigzag persistence algorithms [14, 15]. The folding process is the only costly step for which we provide an efficient algorithm.
One could also argue that a full decomposition algorithm such as the one in [18] or the well-known Meataxe algorithm can be used to compute the generalized rank because they compute all full interval summands. However, the algorithm in [18] does not work for nondistinctly graded modules and the Meataxe algorithm has a high time complexity ($O(t^{18})$, $t$ is the maximum of poset and filtration sizes), as pointed out in [18]). Furthermore, these algorithms expect the input in matrix forms (presentations or linear maps) instead of filtrations that are common in practice. We alleviate these issues; more precisely, highlights of our approach are:
  • It introduces an unfolding/folding technique for $P$-modules, that is, finite dimensional persistence modules defined on a finite poset $P$, which may be of independent interest.
  • It provides an algorithm that, given a simplicial $P$-filtration of a simplicial complex inducing a $P$-module $\mathbb{M}$ by homology functor, computes the generalized rank ${\sf{rk}}(\mathbb{M})$ in $O(t^{\omega+2})$ time, $\omega<2.373$, where the description size of $P$ and the given filtration is at most $t$ (see Eq. (9) and Theorem 6.2). The algorithm does not need to go through an extra step of computing a presentation of $\mathbb{M}$ from its inducing $P$-filtration. In fact, currently all published efficient algorithms for computing presentations work for $2$-parameter modules ($P\subseteq\mathbb{R}^{2},\mathbb{Z}^{2}$[24, 28] and not for a general $P$-module indexed by a finite poset $P$.
  • It computes full interval summands of $\mathbb{M}$ representing its “global sections" supported on $P$.
  • It gives a more efficient $O(t^{\omega})$ algorithm for the special cases of degree $d$-homology for $d$-complexes and a linear time algorithm for degree-$1$ homology in graphs.
2 Preliminaries and idea
2.1 Persistence modules
We consider finite dimensional persistence modules indexed by connected finite posets.
Definition 2.1.
For a poset $P$, let $\leq_{P}$ denote the partial order defining it. We also treat $P$ as a category with every $p\in P$ as its object and $\leq_{P}$ inducing the morphisms between them. Two points $p\leq_{P}q$ in $P$ are called immediate and written $p\rightarrow q$ or $q\leftarrow p$ if and only if there is no $r\in P$, $r\not\in\{p,q\}$, satisfying $p\leq_{P}r\leq_{P}q$. We also write $p\leftrightarrow q$ to denote that either $p\rightarrow q$ or $p\leftarrow q$.
Definition 2.2.
An interval $I$ of a poset $P$ is a non-empty subset $I\subseteq P$ so that (i) $I$ is convex with the partial order of $P$, that is, if $p,q\in I$ and $p\leq_{P}r\leq_{P}q$, then $r\in I$ ; (ii) $I$ is connected, that is, for any $p,q\in I$, there is a sequence $p=p_{0},p_{2},\cdots,p_{m}=q$ of elements of $I$ with $p_{i}\leftrightarrow p_{i+1}$ for $i\in\{0,\cdots,m-1\}$. Assuming $P$ is finite and connected, $P$ is also an interval called the the full interval.
Definition 2.3.
Given a poset $P$, we define a $P$-module to be a functor $\mathbb{M}:P\rightarrow\mathbf{vec}_{\mathbb{F}}$ where $\mathbf{vec}_{\mathbb{F}}$ is the category of finite dimensional vector spaces over a fixed field $\mathbb{F}$ with the morphisms being the linear maps among them. For two points $p\leq_{P}q$, we also write $\mathbb{M}(p\leq_{P}q)$ to denote the morphisms of $\mathbb{M}$.
Definition 2.4.
A $P$-module $\mathbb{M}$ is called indecomposable if there is no direct sum $\mathbb{M}\cong\mathbb{M}_{1}\oplus\mathbb{M}_{2}$ so that both $\mathbb{M}_{1}$ and $\mathbb{M}_{2}$ are nontrivial $P$-modules.
Any (pointwise) finite dimensional $P$-module is a direct sum of indecomposables with local endomorphism ring [13]; see also [4]. Such a decomposition is essentially unique up to automorphism according to Azumaya-Krull-Remak-Schmidt theorem [1]: (Also see [26, Theorem 1.11]).
Theorem 2.1.
Every $P$-module has a unique decomposition up to isomorphism $\mathbb{M}\cong\mathbb{M}_{1}\oplus\cdots\oplus\mathbb{M}_{k}$ where each $\mathbb{M}_{i}$ is an indecomposable module.
For a decomposable module $\mathbb{M}$, there exist submodules $\mathbb{M}_{i}$, $i\in[k]$, of $\mathbb{M}$ with inclusions $j_{i}:\mathbb{M}_{i}\rightarrow\mathbb{M}$ so that $(j_{i}:\mathbb{M}_{i}\rightarrow\mathbb{M})_{i\in[k]}$ make $\mathbb{M}$ the direct sum of $\mathbb{M}_{i}$s written as $\mathbb{M}=\mathbb{M}_{1}\oplus\cdots\oplus\mathbb{M}_{k}$. In light of this, we replace the isomorphism in Theorem 2.1 $\mathbb{M}\cong\mathbb{M}_{1}\oplus\cdots\oplus\mathbb{M}_{k}$ with equality $\mathbb{M}=\mathbb{M}_{1}\oplus\cdots\oplus\mathbb{M}_{k}$ where each $\mathbb{M}_{i}$ is indecomposable and is a submodule of $\mathbb{M}$; see Figure 1. We call such a decomposition an internal direct decomposition or simply direct decomposition denoted $\mathcal{D}:=\mathcal{D}(\mathbb{M})$. We treat $\mathcal{D}$ as the set $\{\mathbb{M}_{i}\}_{i\in[k]}$ of indecomposables if there is no confusion. Notice that the uniqueness of such a decomposition is up to automorphisms of $\mathbb{M}$ (and permutations of $\mathbb{M}_{i}$s). This aspect plays an important role in our algorithm to follow.
For an interval $I$ of $P$, any module $\mathbb{I}^{I}:P\rightarrow\mathbf{vec}_{\mathbb{F}}$ is an interval module if:
$\mathbb{I}^{I}(p)\cong\begin{cases}\mathbb{F}&\mbox{if}\ p\in I,\\ 0&\mbox{otherwise,}\end{cases}\hskip 42.67912pt\mathbb{I}^{I}(p\leq_{P}q)\cong\begin{cases}\mathrm{id}_{\mathbb{F}}&\mbox{if}\,\,p,q\in I,\ p\leq_{P}q,\\ 0&\mbox{otherwise.}\end{cases}$
Definition 2.5.
An interval module $\mathbb{I}=\mathbb{I}^{P}$ $(\mathbb{I}(p)\neq 0$ for any $p\in P)$ is a full interval module(Figure 1).
Figure 1: (left) A $P$-module with vector spaces generated by bases shown at the points of $P$ (arrows represent the partial order) and internal maps shown as matrices; (right) a direct decomposition that contains two full interval modules (top,middle) and another interval module (bottom) which is not full.
2.2 Limits and colimits and generalized rank
We recall the notions of limit and colimit from category theory [29]. Although it is known that limit and colimit may not exist for all functors, they do exist for functors defined on finite posets, which is the case we consider. The following definitions are reproduced from [16]. Let $\mathcal{C}$} denote a category in the following definitions.
Definition 2.6 (Cone).
Let $F:P\rightarrow{\mathcal{C}}$ be a functor. A cone over $F$ is a pair $\left(L,(\pi_{p})_{p\in P}\right)$ consisting of an object $L$ in $\mathcal{C}$} and a collection $(\pi_{p})_{p\in P}$ of morphisms $\pi_{p}:L\rightarrow F(p)$ that commute with the arrows in the diagram of $F$, i.e. if $p\leq q$ in $P$, then $\pi_{q}=F(p\leq q)\circ\pi_{p}$ in $\mathcal{C}$}, i.e. the diagram below commutes.
[…]
The cone $\left(L,(\pi_{p})_{p\in P}\right)$ over $F$ sometimes is denoted simply by $L$, suppressing the morphisms $(\pi_{p})_{p\in P}$ if it causes no confusion. A limit of $F:P\rightarrow{\mathcal{C}}$ is an object in the collection of all cones over $F$ that is terminal:
Definition 2.7 (Limit).
Let $F:P\rightarrow{\mathcal{C}}$ be a functor. A limit of $F$ is a cone over $F$, denoted by $\left(\mathsf{lim}\,F,\ (\pi_{p})_{p\in P}\right)$ or simply $\mathsf{lim}\,F$, with the following (universal) terminal property: For any cone $\left(L^{\prime},(\pi^{\prime}_{p})_{p\in P}\right)$ of $F$, there is a unique morphism $u:L^{\prime}\rightarrow\mathsf{lim}\,F$ such that $\pi_{p}^{\prime}=\pi_{p}\circ u$ for all $p\in P$.
Cocones and colimits are defined in a dual manner:
Definition 2.8 (Cocone).
Let $F:P\rightarrow{\mathcal{C}}$ be a functor. A cocone over $F$ is a pair $\left(C,(i_{p})_{p\in P}\right)$ consisting of an object $C$ in $\mathcal{C}$} and a collection $(i_{p})_{p\in P}$ of morphisms $i_{p}:F(p)\rightarrow C$ that commute with the arrows in the diagram of $F$, i.e. if $p\leq q$ in $P$, then $i_{p}=i_{q}\circ F(p\leq q)$ in $\mathcal{C}$}, i.e. the diagram below commutes.
[…]
In Definition 2.8, a cocone $\left(C,(i_{p})_{p\in P}\right)$ over $F$ is sometimes denoted simply by $C$, suppressing the collection $(i_{p})_{p\in P}$ of morphisms. A colimit of $F:P\rightarrow{\mathcal{C}}$ is an initial object in the collection of cocones over $F$:
Definition 2.9 (Colimit).
Let $F:P\rightarrow{\mathcal{C}}$ be a functor. A colimit of $F$ is a cocone, denoted by $\left(\mathsf{colim}\,F,\ (i_{p})_{p\in P}\right)$ or simply $\mathsf{colim}\,F$, with the following initial property: If there is another cocone $\left(C^{\prime},(i^{\prime}_{p})_{p\in P}\right)$ of $F$, then there is a unique morphism $u:\mathsf{colim}\,F\rightarrow C^{\prime}$ such that $i^{\prime}_{p}=u\circ i_{p}$ for all $p\in P$.
The following proposition gives a standard way of constructing a limit of a $P$-module $\mathbb{M}$. See for example [16, 25]).
Notation 2.1.
Let $p,q\in P$ and let $v_{p}\in\mathbb{M}(p)$ and $v_{q}\in\mathbb{M}(q)$. We write $v_{p}\sim v_{q}$ if $p$ and $q$ are comparable, and either $v_{p}$ is mapped to $v_{q}$ via $\mathbb{M}(p\leq_{P}q)$ or $v_{q}$ is mapped to $v_{p}$ via $\mathbb{M}(q\leq_{P}p)$.
Proposition 2.1.
  1. (i)
    The limit of $\mathbb{M}$ is (isomorphic to) the pair $\left(W,(\pi_{p})_{p\in P}\right)$ described as follows:
    $W:=\left\{(v_{p})_{p\in P}\in\bigoplus_{p\in P}\mathbb{M}(p):\ \forall p\leq q\mbox{ in }P,\ v_{p}\sim v_{q}\right\}$ (3)
    and for each $p\in P$, the map $\pi_{p}:W\rightarrow\mathbb{M}(p)$ is the canonical projection. An element of $W$ is called a global section of $\mathbb{M}$.
  2. (ii)
    The colimit of $\mathbb{M}$ is (isomorphic to) the pair $\left(U,(i_{p})_{p\in P}\right)$ described as follows: For $p\in P$, let the map $j_{p}:\mathbb{M}(p)\hookrightarrow\bigoplus_{p\in P}\mathbb{M}(p)$ be the canonical injection. The space $U$ is the quotient space $\left(\bigoplus_{p\in P}\mathbb{M}(p)\right)/T$, where $T$ is the subspace of $\bigoplus_{p\in P}\mathbb{M}(p)$ which is generated by the vectors of the form $j_{p}(v_{p})-j_{q}(v_{q}),\ v_{p}\sim v_{q},$ the map $i_{p}:\mathbb{M}(p)\rightarrow U$ is the composition $\rho\circ j_{p}$, where $\rho$ is the quotient map $\bigoplus_{p\in P}\mathbb{M}(p)\rightarrow U$.
These definitions imply that, for every $p\leq_{P}q$ in $P$, $\mathbb{M}(p\leq_{P}q)\circ\pi_{p}=\pi_{q}\ \ \mbox{and }\ i_{q}\circ\mathbb{M}(p\leq_{P}q)=i_{p}$, which in turn imply $i_{p}\circ\pi_{p}=i_{q}\circ\pi_{q}:L\rightarrow C$ for any $p,q\in P$.
Definition 2.10 ([25]).
The canonical limit-to-colimit map $\psi_{\mathbb{M}}:\mathsf{lim}\,\mathbb{M}\rightarrow\mathsf{colim}\,\mathbb{M}$ is the linear map $i_{p}\circ\pi_{p}$ for any $p\in P$. The generalized rank of $\mathbb{M}$ is ${\sf{rk}}(\mathbb{M}):=\mathrm{rank}(\psi_{\mathbb{M}})$.
The following result allows us to compute ${\sf{rk}}(\mathbb{M})$ as the number of the full interval modules in a direct decomposition of $\mathbb{M}$.
Theorem 2.2 ([11, Lemma 3.1]).
The rank ${\sf{rk}}(\mathbb{M})$ is equal to the number of full interval modules in a direct decomposition of $\mathbb{M}$.
2.3 Idea of unfolding to zigzag module
The overall idea of our approach is to “straighten up” the given $P$-module $\mathbb{M}$ into a zigzag module $\mathbb{M}_{ZZ}$ which is a zigzag module defined over a linear poset $P_{ZZ}$. It is well known that a zigzag module like $\mathbb{M}_{ZZ}$ decomposes into interval modules and efficient algorithms for computing them exist. After computing these interval modules, we attempt to fold back the full interval modules in this decomposition of $\mathbb{M}_{ZZ}$ to full interval summands of the original module $\mathbb{M}$. We know that full interval modules that are summands of $\mathbb{M}$ unfolds into full interval modules that are summands of $\mathbb{M}_{ZZ}$. However, the converse may not hold, that is, not all full interval summands of $\mathbb{M}_{ZZ}$ fold back to full interval summands of $\mathbb{M}$. So, the main challenge becomes to determine which full interval modules in the computed decomposition of $\mathbb{M}_{ZZ}$ do indeed fold back (possibly with some modifications) to full interval summands of $\mathbb{M}$.
Figure 2: Unfolding two modules (defined over $\mathbb{Z}_{2}$) on the same poset into zigzag modules; hollow circles represent zero vectors. (top row) Decomposition of the zigzag module folds back to two interval modules, one with support on the entire poset giving rise to a generalized rank equal to one, and the other with support on $D$. (bottom row) Full interval module in the decomposition of the zigzag module does not fold back to a full interval module (attempted folding shows the discrepancy at point $D$), giving rise to generalized rank equal to zero.
Figure 2 illustrates this idea. The module $\mathbb{M}$ is defined on a four point poset $P=\{A,B,C,D\}$ shown on left. Bases of the vector spaces are shown in open brackets ‘$()$’ and linear maps in these bases are shown as matrices. The poset is straightened into a zigzag path $A\rightarrow D\rightarrow C\leftarrow D\leftarrow B$. One way to look at this straightening is to view $P$ as a directed graph and to traverse all its vertices and edges sequentially possibly with repetition. Starting from the vertex $A$ and moving to an adjacent node disregarding the direction while noting down the visited node and the directed edge produces the zigzag path. This process of unfolding a poset into a zigzag path is formalized in section 5.1. The module $\mathbb{M}$ is unfolded into the zigzag module $\mathbb{M}_{ZZ}$ by copying the vector spaces and linear maps at vertices and edges respectively into their unfolded versions. For the module shown on the top, we get three interval modules (bars) in a decomposition of $\mathbb{M}_{ZZ}$, the full interval module supported on $[A,B]$ and the other two supported on two copies of $D$ respectively. Bases of one dimensional vector spaces for the interval modules are indicated beneath them. When we fold back $P_{ZZ}$ to $P$ (reversing the process of unfolding) sending $\mathbb{M}_{ZZ}$ to $\mathbb{M}$, the full interval module does fold back to a summand that is a full interval module because the vectors $v_{1}+v_{2}$111this is a vector addition, not a direct sum at two copies of $D$ are the same and hence map to the same vector in $\mathbb{M}(D)$. The other two interval modules supported on single-points in the decomposition of $\mathbb{M}_{ZZ}$ also fold back to a summand of $\mathbb{M}$ supported on a single point $D$ where $\mathbb{M}(D)$ is generated by the vector $v_{2}$.
The case for the module shown in the bottom row of Figure 2 is not the same. In this case, $\mathbb{M}_{ZZ}$ also has the barcode consisting of the same three intervals, but the corresponding interval modules are not the same. The full interval module in this case has different vector spaces spanned by $v_{1}$ and $v_{2}$ respectively at the two copies of $D$. Thus, this interval module does not fold into a full interval submodule of $\mathbb{M}$ as an attempt on right indicates. We can determine such full interval modules in a decomposition of the zigzag module by checking if the vectors at the copied vertices are the same or not. However, even if this check succeeds, the interval module may not fold to a summand of $\mathbb{M}$. Figure 3 (top and bottom) shows such examples. On the other hand, even if this check fails, it may be possible to change the full interval module to have the vectors to be the same at the copied vertices. Figure 4 in section 4.2 illustrates such an example. Determining such cases and taking actions accordingly are key aspects of our algorithm.
3 Folding and Unfolding
As already indicated, our algorithm operates on two main constructs called folding and unfolding of persistence modules. In this section, we introduce these concepts formally for general modules and then later specialize them to zigzag modules for the algorithm.
Definition 3.1.
Let $Q$ be a finite poset. A poset $\mathrm{Fld}_{s}Q$ is a folded poset of $Q$ if there exists a surjection $s:Q\rightarrow\mathrm{Fld}_{s}Q$, which (i) preserves order, that is, $p\leq_{Q}q$ only if $s(p)\leq_{\mathrm{Fld}_{s}Q}s(q)$ for all $p,q\in Q$, and (ii) surjects also on the Hasse diagram of $\mathrm{Fld}_{s}Q$, that is, for every immediate pair $u\rightarrow v$ in $\mathrm{Fld}_{s}Q$, there is a pair $p\leq_{Q}q$ where $s(p)=u$ and $s(q)=v$. We say $s$ is a folding of $Q$ and $Q$ is an unfolded poset of $\mathrm{Fld}_{s}Q$.
Viewing posets as categories, a folding $s:Q\rightarrow\mathrm{Fld}_{s}Q$ can be viewed as a functor from $Q$ to $\mathrm{Fld}_{s}Q$.
Definition 3.2.
Let $P=\mathrm{Fld}_{s}Q$ and $\mathbb{M}:P\to\mathbf{vec}_{\mathbb{F}}$ be a $P$-module and $\mathbb{N}:Q\to\mathbf{vec}_{\mathbb{F}}$ be a $Q$-module. We say $\mathbb{M}$ is an $s$-folding of $\mathbb{N}$ $(\mathbb{N}$ $s$-folds or simply folds into $\mathbb{M})$ and $\mathbb{N}$ is an $s$-unfolding of $\mathbb{M}$ $(\mathbb{M}$ $s$-unfolds or simply unfolds into $\mathbb{N})$ if there is a folding $s:Q\rightarrow P$ so that
$\mathbb{N}(q)=\mathbb{M}(s(q))\,~(equality~as~sets)~\forall q\in Q\,;\,\mathbb{N}(p\leq_{Q}q)=\mathbb{M}(s(p)\leq_{P}s(q))\,\forall(p\leq_{Q}q).$ (4)
We write $\mathbb{M}=\mathrm{Fld}_{s}(\mathbb{N})$ and $\mathbb{N}=\mathrm{Fld}_{s}^{-1}(\mathbb{M})$.
Remark 3.1.
Observe that for a given folding $s:Q\to P$, a $P$-module $\mathbb{M}$ always has an induced $s$-unfolding $\mathbb{M}\circ s$ by pre-composition with $s$. However, for a given $Q$-module $\mathbb{N}$, an $s$-folding may not exist because it may happen that $\mathbb{N}(q)\not=\mathbb{N}(q^{\prime})$ where $s(q^{\prime})=s(q)$, or $\mathbb{N}(q_{1}\rightarrow q_{2})\not=\mathbb{N}(q_{1}^{\prime}\rightarrow q_{2}^{\prime})$ where $s(q_{1})=s(q_{1}^{\prime})$ and $s(q_{2})=s(q_{2}^{\prime})$.
An interesting and important fact is that two isomorphic modules may have different $s$-foldings. Figure 2 shows such an example. Two zigzag modules shown in the middle are isomorphic (barcodes are the same), but they are not exactly the same as modules (even though vector spaces are pointwise equal, morphisms are not). So, even if they are isomorphic, they fold to different modules as shown in left. Nevertheless, if a folding exists, a module necessarily folds to a unique module as Proposition 3.1 states.
Definition 3.3.
Let $\mathbb{M}_{1}$ be a summand of $\mathbb{M}$. A complement of $\mathbb{M}_{1}$ denoted $\overline{\mathbb{M}_{1}}$ is a summand so that $\mathbb{M}=\mathbb{M}_{1}\oplus\overline{\mathbb{M}_{1}}$. Observe that $\overline{\mathbb{M}_{1}}$ is not necessarily unique though for a given decomposition, it is uniquely identified.
Definition 3.4.
For a folding $s:Q\rightarrow\mathrm{Fld}_{s}(Q)$ and a $Q$-module $\mathbb{N}$, we say $\mathbb{N}$ is $s$-foldable ((or simply foldable)) if $\mathbb{N}(q)=\mathbb{N}(q^{\prime})$ for every pair $q,q^{\prime}\in Q$ where $s(q)=s(q^{\prime})$.
We weaken the condition of foldability to invertibility in the following definition.
Definition 3.5.
For a folding $s:Q\rightarrow\mathrm{Fld}_{s}(Q)$ and a $Q$-module $\mathbb{N}$, a summand $\mathbb{N}^{\prime}$ in a decomposition $\mathbb{N}=\mathbb{N}^{\prime}\oplus\mathbb{N}$$^{\prime\prime}$ is called invertible if there exists a foldable summand $\mathbb{Q}$ of $\mathbb{N}$ so that $\mathbb{N}=\mathbb{Q}\oplus\mathbb{N}^{\prime\prime}$.
Foldability of a summand and its complement of a module guarantees that they both remain summands after the folding of the module. However, we can weaken the requirement on foldability for the complement. If foldability holds for a summand then it remains a summand after the folding if and only if its complement is invertible. These are the statements of the following theorem.
Theorem 3.1.
Let $\mathbb{M}$ be a $P$-module and $\mathbb{N}$ be a $Q$-module where $\mathbb{M}=\mathrm{Fld}_{s}(\mathbb{N})$ for some folding $s:Q\rightarrow P$.
  1. 1.
    If $\mathbb{N}=\mathbb{N}_{1}\oplus\overline{\mathbb{N}_{1}}$ and both $\mathbb{N}_{1}$ and $\overline{\mathbb{N}_{1}}$ are foldable, then $\mathrm{Fld}_{s}(\mathbb{N}_{1})$ and $\mathrm{Fld}_{s}(\overline{\mathbb{N}_{1}})$ exist and $\mathbb{M}=\mathrm{Fld}_{s}(\mathbb{N}_{1})\oplus\mathrm{Fld}_{s}(\overline{\mathbb{N}_{1}})$.
  2. 2.
    Conversely, if $\mathbb{M}=\mathbb{M}_{1}\oplus\overline{\mathbb{M}_{1}}$, then $\mathrm{Fld}_{s}^{-1}(\mathbb{M}_{1})$ and $\mathrm{Fld}_{s}^{-1}(\overline{\mathbb{M}_{1}})$ necessarily exist and $\mathbb{N}=\mathrm{Fld}_{s}^{-1}(\mathbb{M}_{1})\oplus\mathrm{Fld}_{s}^{-1}(\overline{\mathbb{M}_{1}})$.
  3. 3.
    If $\mathbb{N}=\mathbb{N}_{1}\oplus\overline{\mathbb{N}_{1}}$ where $\mathbb{N}_{1}$ is foldable, then $\mathrm{Fld}_{s}({\mathbb{N}_{1}})$ necessarily exists and is a summand of $\mathbb{M}$ if and only if $\overline{\mathbb{N}_{1}}$ is invertible.
In Figure 3 (bottom), the gray interval is foldable and its complement (direct sum of the red and blue intervals) is foldable. So, the gray interval and its complement fold into summands whereas in the top figure in Figure 3, none of the interval modules folds into a summand because even if the blue one is foldable, its complement red one is neither foldable nor invertible.
Figure 3: (left) $\mathbb{M}$, (right) $\mathbb{M}_{ZZ}$: (top) the blue interval module is foldable whereas its complement (red) is not invertible, so neither folds to a summand of $\mathbb{M}$; (bottom) the top gray interval module is foldable and its complement is foldable (hence invertible), so the gray interval module folds to a summand of $\mathbb{M}$. The blue interval module is foldable, but its complement is not invertible (foldable), so it does not fold to a summand in $\mathbb{M}$.
The following propositions help proving Theorem 3.1.
Proposition 3.1.
For a $Q$-module $\mathbb{N}$ and a folding $s:Q\to P$, if $\mathrm{Fld}_{s}(\mathbb{N})$ exists, then $\mathrm{Fld}_{s}(\mathbb{N})$ is unique.
Proof.
If there were two modules $\mathbb{M}_{1}$ and $\mathbb{M}_{2}$ that are $s$-foldings of $\mathbb{N}$, then both $\mathbb{M}_{1}(p)=\mathbb{M}_{2}(p)=\mathbb{N}(s^{-1}(p))$ for every $p\in P$. Furthermore, $\mathbb{M}_{1}(p\leq_{P}p^{\prime})=\mathbb{M}_{2}(p\leq_{P}p^{\prime})=\mathbb{N}(s^{-1}(p)\leq_{Q}s^{-1}(p^{\prime}))$ $\forall(p\leq_{P}p^{\prime})$ by definition. This immediately shows that $\mathbb{M}_{1}=\mathbb{M}_{2}$. ∎
Proposition 3.2, Proposition 3.3, and Proposition 3.4 below are used to prove Theorem 3.1 that characterizes modules which fold into summand modules.
Proposition 3.2 ([8]).
Let $\mathbb{N}$ be a submodule of a $P$-module $\mathbb{M}$ where there is a submodule $\overline{\mathbb{N}}$ of $\mathbb{M}$ so that $\mathbb{M}(p)=\mathbb{N}(p)\oplus\overline{\mathbb{N}}(p)$ for every $p\in P$. Then, $\mathbb{N}$ is a summand of $\mathbb{M}$, that is, $\mathbb{M}=\mathbb{N}\oplus\overline{\mathbb{N}}$.
Proposition 3.3.
Let $\mathbb{N}$ be a $Q$-module where $\mathrm{Fld}_{s}(\mathbb{N})$ exists for some folding $s:Q\rightarrow P$. Then, for any submodule $\mathbb{N}^{\prime}\subseteq\mathbb{N}$ that is foldable $\mathrm{Fld}_{s}(\mathbb{N}^{\prime})$ exists.
Proof.
Construct a $P$-module $\mathbb{M}$ as follows: First, put $\mathbb{M}(p)=\mathbb{N}^{\prime}(q)$ where $p=s(q)$. This is well defined because $\mathbb{N}^{\prime}$ is foldable. Next, put $\mathbb{M}(p\leq_{P}p^{\prime})=\mathbb{N}^{\prime}(q\leq_{Q}q^{\prime})$ where $p=s(q)$ and $p^{\prime}=s(q^{\prime})$. This is also well defined because for every pair $q\leq_{Q}q^{\prime}$ so that $p=s(q)$ and $p^{\prime}=s(q^{\prime})$, we have $\mathbb{N}^{\prime}(q\leq_{Q}q^{\prime})$ to be a restriction of $\mathbb{N}(q\leq_{Q}q^{\prime})$ where $\mathrm{Fld}_{s}(\mathbb{N})$ exists. Observe that $\mathbb{M}=\mathrm{Fld}_{s}(\mathbb{N}^{\prime})$ by Definition 3.2. ∎
Proposition 3.4.
Let $\mathbb{M}$ be a $P$-module and $\mathbb{N}$ be a $Q$-module where $\mathbb{M}=\mathrm{Fld}_{s}(\mathbb{N})$ for some folding $s:Q\rightarrow P$. For any submodule $\mathbb{N}^{\prime}\subseteq\mathbb{N}$, if $\mathrm{Fld}_{s}(\mathbb{N}^{\prime})$ exists, then it is a submodule of $\mathbb{M}$. Conversely, if $\mathbb{M}^{\prime}$ is a submodule of $\mathbb{M}$, then $\mathrm{Fld}_{s}^{-1}(\mathbb{M}^{\prime})$ is a submodule of $\mathbb{N}$.
Proof.
We have $\mathbb{N}^{\prime}(q)\subseteq\mathbb{N}(q)$ for $\forall q\in Q$ and $\mathbb{N}^{\prime}(p\leq_{Q}q)$ is a restriction of $\mathbb{N}(p\leq_{Q}q)$ on $\mathbb{N}^{\prime}(p)$ $\forall p\leq_{Q}q$ because $\mathbb{N}^{\prime}$ is a submodule of $\mathbb{N}$. Then, by Definition 3.2,
$\mathrm{Fld}_{s}(\mathbb{N}^{\prime})(s(q))$ $=$ $\mathbb{N}^{\prime}(q)\subseteq\mathbb{N}(q)=\mathbb{M}(s(q))~~\forall q\in Q\mbox{ and }$
$\mathrm{Fld}_{s}(\mathbb{N}^{\prime})(s(p)\leq_{P}s(q))$ $=$ $\mathbb{N}^{\prime}(p\leq_{Q}q)=\mathbb{N}(p\leq_{Q}q)|_{\mathbb{N}^{\prime}(p)}=\mathbb{M}(s(p)\leq_{P}s(q))|_{\mathrm{Fld}_{s}(\mathbb{N}^{\prime})(s(p))},$
which establishes that $\mathrm{Fld}_{s}(\mathbb{N}^{\prime})$ is a submodule of $\mathbb{M}$.
For the converse statement, check that $\mathrm{Fld}_{s}^{-1}(\mathbb{M}^{\prime})$ necessarily exists and it is a submdoule of $\mathbb{M}$ by definition of unfolding. ∎
Proof of Theorem 3.1.
(1) By Proposition 3.3, both $\mathrm{Fld}_{s}(\mathbb{N}_{1})$ and $\mathrm{Fld}_{s}(\overline{\mathbb{N}_{1}})$ exist and they are submodules of $\mathbb{M}$ by Proposition 3.4. If we show that $\mathbb{M}(p)=\mathrm{Fld}_{s}(\mathbb{N}_{1})(p)\oplus\mathrm{Fld}_{s}(\overline{\mathbb{N}_{1}})(p)$ for every $p\in P$, then $\mathbb{M}=\mathrm{Fld}_{s}(\mathbb{N}_{1})\oplus\mathrm{Fld}_{s}(\overline{\mathbb{N}_{1}})$ due to Proposition 3.2.
We have $\mathbb{M}(p)=\mathrm{Fld}_{s}(\mathbb{N})(p)=\mathrm{Fld}_{s}(\mathbb{N}_{1}\oplus\overline{\mathbb{N}_{1}})(p)=\mathrm{Fld}_{s}(\mathbb{N}_{1})(p)\oplus\mathrm{Fld}_{s}(\overline{\mathbb{N}_{1}})(p)$.
(2) For the converse, observe that $\mathrm{Fld}_{s}^{-1}(\mathbb{M}_{1})$ and $\mathrm{Fld}_{s}^{-1}(\overline{\mathbb{M}_{1}})$ necessarily exist and they are submodules of $\mathbb{N}$ by Proposition 3.4. Furthermore, $\mathbb{N}(q)=\mathrm{Fld}_{s}^{-1}(\mathbb{M})(q)=(\mathrm{Fld}_{s}^{-1}(\mathbb{M}_{1}\oplus\overline{\mathbb{M}_{1}}))(q)=\mathrm{Fld}_{s}^{-1}(\mathbb{M}_{1})(q)\oplus\mathrm{Fld}_{s}^{-1}(\overline{\mathbb{M}_{1}})(q)$. Then, by Proposition 3.2, we have $\mathbb{N}=\mathrm{Fld}_{s}^{-1}(\mathbb{M}_{1})\oplus\mathrm{Fld}_{s}^{-1}(\overline{\mathbb{M}_{1}})$.
(3) First, assume that $\mathrm{Fld}_{s}(\mathbb{N}_{1})$ exists and is a summand of $\mathbb{M}$. Then, we have $\mathbb{M}=\mathrm{Fld}_{s}\mathbb{N}_{1}\oplus\overline{\mathrm{Fld}_{s}\mathbb{N}_{1}}$. Applying (2) above, we have $\mathbb{N}=\mathrm{Fld}_{s}^{-1}(\mathrm{Fld}_{s}\mathbb{N}_{1})\oplus\mathrm{Fld}_{s}^{-1}(\overline{\mathrm{Fld}_{s}\mathbb{N}_{1}})=\mathbb{N}_{1}\oplus\mathrm{Fld}_{s}^{-1}(\overline{\mathrm{Fld}_{s}\mathbb{N}_{1}})$. Let $\mathbb{Q}=\mathrm{Fld}_{s}^{-1}(\overline{\mathrm{Fld}_{s}\mathbb{N}_{1}})$ which is foldable. Then,
$\mathbb{N}_{1}\oplus\overline{\mathbb{N}_{1}}=\mathbb{N}=\mathbb{N}_{1}\oplus\mathbb{Q}$ (5)
Since $\mathbb{Q}$ is foldable, we have that $\overline{\mathbb{N}_{1}}$ is invertible.
Now, assume that $\overline{\mathbb{N}_{1}}$ is invertible. By definition, $\mathbb{N}=\mathbb{N}_{1}\oplus\mathbb{Q}$ for some foldable summand $\mathbb{Q}$. Apply (1) above to conclude that $\mathrm{Fld}_{s}(\mathbb{N}_{1})$ exists and is a summand of $\mathbb{M}$.
4 Limit and complete modules
4.1 Limit modules
We introduce some special types of submodules of a module $\mathbb{M}$ which we call limit modules. These limit modules when further specialized to zigzag modules exhibit additional properties, which we utilize in our algorithm.
Definition 4.1 (Limit module).
We call a submodule $\mathbb{I}\subseteq\mathbb{M}$ a limit module for $\mathbb{M}:P\rightarrow\mathbf{vec}_{\mathbb{F}}$ if (i) $\mathbb{I}=\oplus_{i}\mathbb{I}^{I_{i}}$ where each $\mathbb{I}^{I_{i}}$ is an interval module with $I_{i}\cap I_{j}=\emptyset$ for $i\neq j$ and (ii) for every pair $p\leq_{P}q$, the morphism $\mathbb{I}(p\leq_{P}q)$ is a surjection (this allows some particular types of interval modules to be limit modules, see e.g. Eq.(7)). We call an ordered sequence of vectors $v_{p_{0}},\ldots,v_{p_{m}}$, $v_{p_{i}}\in\mathbb{M}(p_{i})$, a limit representative (or simply representative) if there is a limit module $\mathbb{I}$ for $\mathbb{M}$ so that $\mathbb{I}(p_{i})$ is spanned by the vector $v_{p_{i}}$ for each $i\in\{0,\ldots,m\}$. Conversely, every limit module $\mathbb{I}$ admits a representative ${\sf b}^{\mathbb{I}}_{p_{0}},\ldots,{\sf b}^{\mathbb{I}}_{p_{m}}$ where ${\sf b}^{\mathbb{I}}(p_{i})$ spans $\mathbb{I}(p_{i})$.
In what follows we often say a limit module for $\mathbb{M}$ simply a limit module when $\mathbb{M}$ is understood from the context. The reader can observe that limit representatives are elements of $\mathsf{lim}\,\mathbb{M}$.
The following observations about limit modules help understand the roles they play in the rest of the paper.
First, observe that a limit module $\mathbb{I}$, in general, can be either a full interval module or a direct sum of one or more non-overlapping interval modules such as the ones separated by the red arrows in (7).
Second, observe that some of the interval modules in a direct decomposition of $\mathbb{M}_{ZZ}=\bigoplus_{i}\mathbb{I}_{i}$ may be limit modules.
Third, if $v_{p_{0}},\ldots,v_{p_{m}}$ is a representative of a limit module $\mathbb{I}$, then $\alpha v_{p_{0}},\ldots,\alpha v_{p_{m}}$ is also a representative of $\mathbb{I}$ for any scalar $0\neq\alpha\in\mathbb{F}$. In regard to this fact, we introduce the following notation.
For a limit module $\mathbb{I}\subseteq\mathbb{M}_{ZZ}$, if ${\sf b}^{\mathbb{I}}:{\sf b}^{\mathbb{I}}_{p_{0}},\ldots,{\sf b}^{\mathbb{I}}_{p_{m}}$ is a chosen representative for $\mathbb{I}$, then for $0\not=\alpha\in\mathbb{F}$, $\alpha{\sf b}^{\mathbb{I}}$ denotes the representative $\alpha{\sf b}^{\mathbb{I}}_{p_{0}},\ldots,\alpha{\sf b}^{\mathbb{I}}_{p_{m}}$.
The reader may realize that a chosen representative of a limit module $\mathbb{I}$ is an element in $\mathsf{lim}\,\mathbb{I}$ representing a global section of $\mathbb{M}_{ZZ}$. They can be added to produce other sections. This addition is given by pointwise vector addition which should not be confused with direct sums.
Observation 4.1 (representative sums).
For two limit modules $\mathbb{I}$ and $\mathbb{I}^{\prime}$ and for $\alpha,\alpha^{\prime}\in\mathbb{F}$, the sequence of vectors $(\alpha{\sf b}^{\mathbb{I}}_{p_{0}}+\alpha^{\prime}{\sf b}^{\mathbb{I}^{\prime}}_{p_{0}}),\ldots,(\alpha{\sf b}^{\mathbb{I}}_{p_{m}}+\alpha^{\prime}{\sf b}^{\mathbb{I}^{\prime}}_{p_{m}})$ is a representative. We denote this representative as the sum $\alpha{\sf b}^{\mathbb{I}}+\alpha^{\prime}{\sf b}^{\mathbb{I}^{\prime}}$.
The representative $\alpha{\sf b}^{\mathbb{I}}+\alpha^{\prime}{\sf b}^{\mathbb{I}^{\prime}}$ can be viewed as an element in the space $\mathsf{lim}\,\mathbb{I}\oplus\mathsf{lim}\,\mathbb{I}^{\prime}$ obtained by fixing an element ${\sf b}^{\mathbb{I}}$ in $\mathsf{lim}\,\mathbb{I}$ and ${\sf b}^{\mathbb{I}^{\prime}}$ in $\mathsf{lim}\,\mathbb{I}^{\prime}$ and mapping them to the direct sum by inclusions.
4.2 Complete modules
Definition 4.2.
Let $P=\mathrm{Fld}_{s}Q$ and $\mathbb{N}:Q\rightarrow\mathbf{vec}_{\mathbb{F}}$ be a $Q$-module so that $\mathbb{M}=\mathrm{Fld}_{s}(\mathbb{N})$ exists. An interval module $\mathbb{I}$ in a direct decomposition $\mathcal{D}:\bigoplus_{i}\mathbb{I}_{i}$ of $\mathbb{N}$ is called $s$-complete if and only if (i) $\mathbb{I}$ is a full interval module that is foldable in $\mathcal{D}$ and (ii) its complement $\overline{\mathbb{I}}$ in $\mathcal{D}$ is invertible. Let $\kappa(\mathcal{D})$ denote the number of $s$-complete interval modules in $\mathcal{D}$. We call $\mathcal{D}$ $s$-complete if $\kappa(\mathcal{D})={\sf{rk}}(\mathbb{M})$.
Theorem 3.1 helps us to prove the following proposition which guarantees that a foldable module $\mathbb{N}$ has a $s$-complete decomposition. Additionally, it states that no direct decomposition of $\mathbb{N}$ can have more $s$-complete interval modules than a $s$-complete decomposition can have.
Proposition 4.1.
Let $P=\mathrm{Fld}_{s}Q$ be a folded poset of a finite poset $Q$. Let $\mathbb{N}:Q\rightarrow\mathbf{vec}_{\mathbb{F}}$ be a $Q$-module so that the $s$-folding $\mathbb{M}=\mathrm{Fld}_{s}(\mathbb{N})$ exists. Then, a $s$-complete decomposition of $\mathbb{N}$ exists. Furthermore, any direct decomposition $\mathcal{D}$ of $\mathbb{N}$ has $\kappa(\mathcal{D})\leq{\sf{rk}}(\mathbb{M})$.
Proof.
First, we prove the second conclusion. If a direct decomposition $\mathcal{D}$ of $\mathbb{N}$ had $\kappa(\mathcal{D})>{\sf{rk}}(\mathbb{M})$, then $\mathcal{D}$ would have more than ${\sf{rk}}(\mathbb{M})$ $s$-complete interval modules as its summand each of which would fold to a full interval summand of $\mathbb{M}$ (Theorem 3.1(3)). This is not possible because in that case $\mathbb{M}$ would have more than ${\sf{rk}}(\mathbb{M})$ summands that are full intervals, an impossibility according to Theorem 2.2.
Next, we show the first conclusion. Consider a direct decomposition $\mathbb{M}=\mathbb{I}_{1}\oplus\cdots\oplus\mathbb{I}_{r}\oplus\mathbb{M}^{\prime}_{1}\oplus\cdots\oplus\mathbb{M}^{\prime}_{k}$ where $\mathbb{I}_{1},\ldots,\mathbb{I}_{r}$ are full interval modules. By Theorem 2.2, $r={\sf{rk}}(\mathbb{M})$. Then, there is a decomposition $\mathcal{D}:\mathrm{Fld}_{s}^{-1}(\mathbb{I}_{1})\oplus\cdots\oplus\mathrm{Fld}_{s}^{-1}(\mathbb{I}_{r})\oplus\mathrm{Fld}_{s}^{-1}(\mathbb{M}^{\prime}_{1})\oplus\cdots\oplus\mathrm{Fld}_{s}^{-1}(\mathbb{M}^{\prime}_{k})$ of $\mathbb{N}$ by Theorem 3.1(2). Furthermore, each of $\mathrm{Fld}_{s}^{-1}(\mathbb{I}_{i})$, $1\leq i\leq r$, is a full interval module because each $\mathbb{I}_{i}$ is so. By definition, both $\mathrm{Fld}_{s}^{-1}(\mathbb{I}_{i})$ and its complement are foldable (hence also invertible). Therefore, each $\mathrm{Fld}_{s}^{-1}(\mathbb{I}_{i})$ is $s$-complete. A direct decomposition $\mathcal{D}^{\prime}$ (into indecomposables) that refines $\mathcal{D}$ is an $s$-complete decomposition of $\mathbb{N}$ because it has $r={\sf{rk}}(\mathbb{M})$ $s$-complete interval modules and it cannot have any more $s$-complete interval modules as $\kappa(\mathcal{D})\leq{\sf{rk}}(\mathbb{M})$. ∎
5 Zigzag module
Definition 5.1.
A poset $P$ is called a zigzag poset iff there is a linear ordering $p_{0},\ldots,p_{m}$ of the points in $P$, called the zigzag path, so that for $i\in\{0,1,\ldots,m-1\}$, $p_{i}\leftrightarrow p_{i+1}$ are the only and all immediate pairs in $P$, i.e., zigzag path represents the Hasse diagram of $P$. We write $p_{i},p_{j}]$ to denote an interval $I\subseteq P$ with the zigzag path $p_{i},p_{i+1},\cdots,p_{j}$.
Definition 5.2.
A zigzag module $\mathbb{M}_{ZZ}:P_{ZZ}\rightarrow\mathbf{vec}_{\mathbb{F}}$ is a persistence module where the poset $P_{ZZ}$ is a zigzag poset. Assuming that $p_{0},p_{1},\ldots,p_{m}$ is the zigzag path for $P_{ZZ}$, we write the zigzag module as:
$\mathbb{M}_{ZZ}:V_{p_{0}}\stackrel{{\scriptstyle\phi_{0}}}{{\longleftrightarrow}}\ldots\stackrel{{\scriptstyle\phi_{i-1}}}{{\longleftrightarrow}}V_{p_{i}}\stackrel{{\scriptstyle\phi_{i}}}{{\longleftrightarrow}}V_{p_{i+1}}\stackrel{{\scriptstyle\phi_{i+1}}}{{\longleftrightarrow}}\ldots\stackrel{{\scriptstyle\phi_{m-1}}}{{\longleftrightarrow}}V_{p_{m}}$ (6)
where for $i\in\{0,1,\ldots,m-1\}$, $V_{p_{i}}=\mathbb{M}_{ZZ}(p_{i})$ denote the vector spaces and $\phi_{i}=\mathbb{M}_{ZZ}(p_{i}\leq_{P_{ZZ}}p_{i+1})$ or $\mathbb{M}_{ZZ}(p_{i+1}\leq_{P_{ZZ}}p_{i})$ denote the morphisms (linear maps).
We will be interested in interval submodules $\mathbb{I}^{[b_{i},d_{i}]}$ of $\mathbb{M}_{ZZ}$. Such an interval module is either full or can be of four types determined by the types of its end points. The point $b_{i}$ is called closed if $b_{i}\not=0$ and the arrow between $b_{i-1},b_{i}$ is a forward arrow $\rightarrow$’ and is called open otherwise. Similarly, the point $d_{i}$ is called closed if $d_{i}\not=m$ and the arrow between $d_{i},d_{i+1}$ is a backward arrow and is called open otherwise. The interval module $\mathbb{I}^{[b_{i},d_{i}]}$ is called open-open, open-closed, closed-open, or closed-closed depending on whether $b_{i}$ and $d_{i}$ are both open, $b_{i}$ is open and $d_{i}$ closed, $b_{i}$ is closed and $d_{i}$ is open, or both $b_{i}$ and $d_{i}$ are closed respectively. Notice that a full interval module is considered open-open by definition.
By Theorem 2.1, a zigzag module $\mathbb{M}_{ZZ}:P_{ZZ}\rightarrow\mathbf{vec}_{\mathbb{F}}$ over a finite zigzag poset $P_{ZZ}$ also decomposes uniquely into indecomposables. By quiver theory [21, 33], these indecomposables are interval modules, that is, $\mathbb{M}_{ZZ}=\mathbb{I}_{1}\oplus\cdots\oplus\mathbb{I}_{k}$ where each $\mathbb{I}_{i}:=\mathbb{I}^{I_{i}}$ is an interval module defined over an interval $I_{i}:=[p_{b_{i}},p_{d_{i}}]$.
We will be interested in limit modules in direct decompositions of zigzag modules. Observe that limit modules for a zigzag module $\mathbb{M}_{ZZ}$ are dierct sum of open-open interval modules and take the following form.
$0\displaystyle 0\leftrightarrow\cdots{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{0\leftarrow}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{I}({p_{b_{i}}})\leftrightarrow\cdots\leftrightarrow\mathbb{I}({p_{d_{i}}})}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rightarrow 0}\leftrightarrow\cdots\leftrightarrow{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}0\leftarrow}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{I}({p_{b_{j}}})\leftrightarrow\cdots\leftrightarrow\mathbb{I}({p_{d_{j}}})}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\rightarrow 0}\leftrightarrow\cdots\leftrightarrow$ 0 (7)
The following proposition says that in a sense any representative of a full interval submodule of $\mathbb{M}_{ZZ}$ is in the span of the representatives of the limit modules present in any direct decomposition of $\mathbb{M}_{ZZ}$. This property does not extend beyond zigzag module. The persistence module $\mathbb{M}$ shown in Figure 3(top left) has a submodule which is a full interval module. However, it cannot be expressed as a direct sum of any limit modules in a direct decomposition of $\mathbb{M}$ because $\mathbb{M}$ itself is indecomposable and is not even an interval module. Also, see Conclusions (section 8) for further comments on this aspect.
Proposition 5.1.
Let $L$ be the set of limit modules in a direct decomposition $\mathcal{D}$ of $\mathbb{M}_{ZZ}$. For any full interval module $\mathbb{I}\subseteq\mathbb{M}_{ZZ}$, there exist unique $\alpha_{i}\in\mathbb{F}$ so that ${\sf b}^{\mathbb{I}}=\sum_{\mathbb{I}_{i}\in L}\alpha_{i}{\sf b}^{\mathbb{I}_{i}}$ where the sum is defined as in Observation 4.1.
Proof.
Let $\mathbb{I}_{1},\ldots,\mathbb{I}_{k}$ be the set of all interval modules in the direct decomposition $\mathcal{D}$ of the $P_{ZZ}$-module $\mathbb{M}_{ZZ}$. For any $p\in P_{ZZ}$, since $\mathbb{M}_{ZZ}(p)=\text{span}({\sf b}^{\mathbb{I}_{1}}_{p},\cdots,{\sf b}^{\mathbb{I}_{k}}_{p})$, there exist uniquely determined $\alpha^{p}_{i}\in\mathbb{F}$ so that ${\sf b}^{\mathbb{I}}_{p}=\sum_{i}\alpha^{p}_{i}{\sf b}^{\mathbb{I}_{i}}_{p}$ where $\alpha_{i}^{p}$ is taken to be zero if ${\sf b}^{\mathbb{I}_{i}}_{p}=0$. Let $L_{p}=\{\mathbb{I}_{i}\,|\,\alpha^{p}_{i}\not=0\}$ and $L^{\prime}=\cup_{p}L_{p}$. It is not difficult to show that $\alpha^{p}_{i}=\alpha^{q}_{i}:=\alpha_{i}$ for any two points $p,q$ in the support of $\mathbb{I}_{i}$. Then, we can write $L^{\prime}=\{\mathbb{I}_{i}\,|\,\alpha_{i}\not=0\}$. We claim that $L^{\prime}$ is a subset of limit modules (open-open) in $\mathcal{D}$, that is, $L^{\prime}\subseteq L$. If not, there is an interval module $\mathbb{I}^{\prime}\in L^{\prime}$ that is not a limit module, that is, $\mathbb{I}^{\prime}$ has an end point $p_{j}\not\in\{p_{0},p_{m}\}$ satisfying either of the following two cases: (i) the arrow for $\phi_{j-1}$ is forward and the cokernel of $\phi_{j-1}$ restricted to $\mathbb{I}^{\prime}(p_{j})$ is non-zero. It follows that the cokernel of $\phi_{j-1}$ restricted to $\mathbb{I}(p_{j})$ is also non-zero. This is impossible as $\mathbb{I}$ is a full interval module and $p_{j}\not\in\{p_{0},p_{m}\}$, (ii) the arrow for $\phi_{j}$ is backward: again, we reach an impossibility with a similar argument. So, $L^{\prime}\subseteq L$ and it follows that ${\sf b}^{\mathbb{I}}=\sum_{\mathbb{I}_{i}\in L}\alpha_{i}{\sf b}^{\mathbb{I}_{i}}$ where $\alpha_{i}=0$ for $\mathbb{I}_{i}\in L\setminus L^{\prime}$ establishing the claim of the proposition. ∎
Proposition 5.1 suggests the following approach to compute the generalized rank of a given $P$-module $\mathbb{M}$: first unfold $P$ into a zigzag poset (path) $P_{ZZ}$ and construct a zigzag module $\mathbb{M}_{ZZ}$ that is an $s$-unfolding of $\mathbb{M}$. It follows that $\mathrm{Fld}_{s}(\mathbb{M}_{ZZ})$ exists and $\mathbb{M}=\mathrm{Fld}_{s}(\mathbb{M}_{ZZ})$ as required by Proposition 5.1. Then, after computing a direct decomposition of $\mathbb{M}_{ZZ}$ with a zigzag persistence algorithm, convert it to an $s$-complete decomposition and determine how many full interval modules (if any) in this decomposition are $s$-complete.
Figure 4: The full red interval module in $\mathbb{M}_{ZZ}$ as shown in top-middle does not fold to a full interval submodule of $\mathbb{M}$ because the vectors at $D$ mismatch (top-right), but after adding the blue interval module (limit module), it is converted to a full interval module as shown in bottom-middle that folds back to a full interval module (pink) that is a summand of $\mathbb{M}$ (bottom-right).
Consider the $P$-module $\mathbb{M}$ shown in Figure 4 (left). After unfolding the module to a zigzag module (middle), suppose we get a decomposition into interval modules as indicated in the middle-top picture. Just like the example in Figure 2 (bottom), the full interval module in this decomposition of $\mathbb{M}_{ZZ}$ does not fold into a full interval submodule of $\mathbb{M}$ because the representative vectors at the two copies of $D$ do not match. In the example of Figure 2, we could not repair this deficiency. However, now we can do so using the limit modules. Observe that the open-open interval module (blue) supported on $D\rightarrow C\leftarrow D$ is a limit module. We can add its representative to the representative of the full interval module to obtain a new full interval module shown in middle-bottom picture. This new module is complete because it and its complement are foldable and thus the module folds into a full interval summand of $\mathbb{M}$ (Theorem 3.1(1)). Observe that any of the other two limit modules (grey) could also serve the purpose. The algorithm GenRank in section 6 essentially determines whether a full interval module $\mathbb{I}$ in the current decomposition of $\mathbb{M}_{ZZ}$ is foldable, and if not, whether it can be converted to one by adding the chosen representatives of a set of limit modules. This is followed up by a check that determines if the complement of the module $\mathbb{I}$ is invertible. If so, it marks converted $\mathbb{I}$ to be complete.
It is instructive to point out that the blue module which is not a full module can also be made foldable by adding the representative of one of the grey modules, say the left one, to its representative, which has been done to obtain the decomposition shown on bottom right in Figure 4. Our algorithm does not do this because we are interested only on folding full interval modules.
5.1 Unfolding to a zigzag path and zigzag module
A finite poset $P$ is represented by a directed acyclic graph $G=(P,E(P))$ where (i) every directed edge $(p,q)\in E(P)$ satisfies $p\leq_{P}q$ and (ii) every immediate pair $p\rightarrow q$ in $P$ must correspond to an edge $(p,q)\in E(P)$. The size of the poset $|P|$ with such a representation is measured as the total number of vertices and edges in $G$. Given a directed graph $G=(P,E(P))$ for a finite poset $P$, we construct a zigzag poset $P_{ZZ}$ using the concept of Eulerian tour in graphs so that $G_{ZZ}=(P_{ZZ},E(P_{ZZ}))$ represents the zigzag path for $P_{ZZ}$ and $P=\mathrm{Fld}_{s}P_{ZZ}$ for some folding $s$.
Figure 5: Unfolding $P$ (left) to a zigzag poset $P_{ZZ}$; wrapping a thread around (right) gives a zigzag path (bottom) in $P_{ZZ}$
Given a connected graph $G=(V,E)$, an Eulerian tour in $G$ is an ordered sequence of its vertices, possibly with repetitions, $u_{0},\ldots,u_{i},u_{i+1},\ldots u_{t}=u_{0}$ so that every edge $(p,q)\in E$ appears exactly once as a consecutive pair $(p=u_{i},u_{i+1}=q)$ of vertices in the sequence. It is known that if $G$ has even degree at every vertex, then $G$ necessarily has an Eulerian tour which can be computed in $O(|V|+|E|)$ time. We consider the undirected version $\overline{G}=(P,E(P))$ of the poset graph $G=(P,E(P))$ and straighten it up using an Eulerian tour, see Figure 5. However, the graph $\overline{G}$ may not satisfy the vertex degree requirement. So, we double every edge, that is, put a parallel edge in the graph for every edge (this is equivalent to wrapping a thread around as shown in Figure 5). The modified graph $\overline{G}$ then has only even-degree vertices. We compute an Eulerian tour $T$ in $\overline{G}$ and for every adjacent pair of vertices $p,q$ in the tour representing an edge $(p,q)$ in $\overline{G}$ we impose the order $p\leq_{T}q$ if and only if the directed edge $(p,q)$ is in $G$. The poset $(T,\leq_{T})$ is taken as the zigzag poset $P_{ZZ}$ and the tour (zigzag path) as its representation. Clearly, the number of edges in the tour (immediate pairs in $P_{ZZ}$) is at most twice the number of edges in $G$ and the number of vertices in $P_{ZZ}$ is one more than that number. We have
Fact 5.1.
$P=\mathrm{Fld}_{s}P_{ZZ}$ for some folding $s$ where $|P_{ZZ}|\leq 2|P|+1$.
In our algorithm, we assume that $\mathbb{M}$ is implicitly given by a $P$-filtration: A $P$-filtration $F(K)$ of a simplicial complex $K$ is a family of subcomplexes $F(K)=\{K_{p}\subseteq K\}_{p\in P}$ so that $K_{p}\subseteq K_{q}$ if $p\leq_{P}q$. We assume that both $K$ and $P$ are finite. Applying the homology functor $H_{k}(\cdot)$ to the filtration $F(K)$, one obtains a module $\mathbb{M}:=\mathbb{M}_{F(K)}$ in degree $k$ where $\mathbb{M}(p)=H_{k}(K_{p})$ and $\mathbb{M}(p\leq_{P}q):H_{k}(K_{p})\rightarrow H_{k}(K_{q})$ induced by the inclusion $K_{p}\subseteq K_{q}$.
First, we unfold the poset $P$ into a zigzag poset $P_{ZZ}$ using the method described before. Let $s$ be the resulting folding given by Fact 5.1. To unfold $\mathbb{M}:P\rightarrow\mathbf{vec}_{\mathbb{F}}$ into a zigzag module $\mathbb{M}_{ZZ}:P_{ZZ}\rightarrow\mathbf{vec}_{\mathbb{F}}$, we build a zigzag filtration $F_{ZZ}=\{K_{p}\}_{p\in P_{ZZ}}$ by assigning $K_{p}:=K_{s(p)}$. To check that $F_{ZZ}$ is indeed a zigzag filtration, observe that $K_{p}\subseteq K_{q}$ for every $p\leq_{P_{ZZ}}q$ because (i) $s(p)\leq_{P}s(q)$ by definition of folding $s$ and (ii) $K_{s(p)}\subseteq K_{s(q)}$ by definition of the filtration $F(K)$. It can be easily verified that applying the homology functor on $F_{ZZ}$, we get the $s$-unfolding $\mathbb{M}_{ZZ}:P_{ZZ}\rightarrow\mathbf{vec}_{\mathbb{F}}$ of $\mathbb{M}$.
6 Algorithm
The algorithm (GenRank in pseudocode) takes a $P$-fitration $F$ and a degree $k$ for the homology group. First, it $s$-unfolds $P$ to a zigzag path of $P_{ZZ}$ and computes the filtration $F_{ZZ}$ (Step 1). Let $\mathbb{M}$ and $\mathbb{M}_{ZZ}$ be the modules obtained by applying the homology functor in degree $k$ on $F$ and $F_{ZZ}$ respectively as described above. We need to compute interval modules from $F_{ZZ}$ that represents a direct decomposition of $\mathbb{M}_{ZZ}$, i.e., we need a zigzag persistence algorithm that computes the intervals in the barcode with a representative (step 3). A sequence of $k$-cycles $z_{b},\ldots,z_{d}$ constitutes a representative of a limit module $\mathbb{I}:=\mathbb{I}^{[p_{b},p_{d}]}$ if $z_{i}]$ is chosen as ${\sf b}^{\mathbb{I}}_{p_{i}}$ for $p_{i}\in[p_{b},p_{d}]$. The zigzag persistence algorithm in [15] can compute these representatives efficiently. Next, the algorithm checks how many interval modules in the computed decomposition of $\mathbb{M}_{ZZ}$ can be converted to $s$-complete modules which provides ${\sf{rk}}(\mathbb{M})$ according to the definition of $s$-completeness.
Next proposition tells us that it is sufficient to check only the full interval modules in a direct decomposition of $\mathbb{M}_{ZZ}$ if they can be converted to $s$-complete modules.
Proposition 6.1.
Let $\mathbb{M}_{ZZ}$ be an $s$-unfolding of $\mathbb{M}$ and $\mathbb{I}_{1},\ldots,\mathbb{I}_{\ell}$ be the set of limit modules in a direct decomposition $\mathcal{D}$ of $\mathbb{M}_{ZZ}$. There exist unique $\alpha_{i}\in\mathbb{F}$, $i\in[\ell]$, so that every $s$-complete interval module $\mathbb{I}$ in any direct decomposition $\mathcal{D}^{\prime}$ of $\mathbb{M}_{ZZ}$ satisfies that ${\sf b}^{\mathbb{I}}=\sum_{i=1}^{\ell}\alpha_{i}{\sf b}^{\mathbb{I}_{i}}$ where for at least one $i\in[\ell]$, $\mathbb{I}_{i}$ is a full interval module and $\alpha_{i}\not=0$.
Proof.
First, Proposition 5.1 allows us to write
${\sf b}^{\mathbb{I}}=\sum_{i=1}^{\ell}\alpha_{i}{\sf b}^{\mathbb{I}_{i}}\mbox{ for some unique }\alpha_{i}\in\mathbb{F},i\in[\ell].$
Suppose that the claim of the proposition is not true. Fix a point $p\in P_{ZZ}$. We have ${\sf b}^{\mathbb{I}}_{p}=\sum_{i=1}^{\ell}\alpha_{i}{\sf b}^{\mathbb{I}_{i}}_{p}$. Recall the quotient map $\rho$ for colimit in Proposition 2.1(ii). For each vector $v_{i}={\sf b}^{\mathbb{I}_{i}}_{p}$, the quotient vector $\rho(v_{i})$ is a zero element in the colimit $\mathsf{colim}\,\mathbb{M}_{ZZ}$ because the limit module $\mathbb{I}_{i}$ which is not full either has a sequence of vectors $v_{i}\leftrightarrow\cdots\rightarrow 0$ or $0\leftarrow\cdots\leftrightarrow v_{i}$ in its representative and thus $v_{i}\sim 0$ (Notation 2.1). It follows that any representative of $\mathbb{I}$ is sent to a zero element in $\mathsf{colim}\,\mathbb{M}_{ZZ}$. Since $\mathbb{I}$ is $s$-complete, $\mathrm{Fld}_{s}(\mathbb{I})$ exists and its representative is an element of $\mathsf{lim}\,\mathbb{M}$. The definition of the folding implies that the limit-to-colimit map sends the representative of $\mathrm{Fld}_{s}(\mathbb{I})$ also to a zero element in $\mathsf{colim}\,\mathbb{M}$.
Since $\mathbb{I}$ is $s$-complete, it is foldable and its complement is also foldable. Then, $\mathrm{Fld}_{s}(\mathbb{I})$ is a full interval summand of $\mathbb{M}$ according to Theorem 3.1(1). Since the full interval summand $\mathrm{Fld}_{s}(\mathbb{I})$ is sent to zero by the limit-to-colimit map, we have ${\sf{rk}}(\mathbb{M}^{\prime})={\sf{rk}}(\mathbb{M})$ where $\mathbb{M}=\mathrm{Fld}_{s}(\mathbb{I})\oplus\mathbb{M}^{\prime}$. However, ${\sf{rk}}(\mathbb{M}^{\prime})\leq{\sf{rk}}(\mathbb{M})-1$ according to Theorem 2.2 reaching a contradiction. ∎
The above proposition suggests that we try to convert every full interval module $\mathbb{I}$ in a direct decomposition $\mathcal{D}$ of $\mathbb{M}_{ZZ}$ to a foldable module first, and then appealing to Theorem 3.1(3) check if the complement module $\overline{\mathbb{I}}$ in $\mathcal{D}$ is invertible.
Definition 6.1.
Let $\mathbb{M}_{ZZ}$ be an $s$-unfolding of $\mathbb{M}$. We say a full interval module $\mathbb{I}$ in a direct decomposition $\mathcal{D}$ of $\mathbb{M}_{ZZ}$ is convertible in $\mathcal{D}$ if either (i) $\mathbb{I}$ is foldable, or (ii) there exists a set of limit modules $\{\mathbb{I}_{i}\}$ in $\mathcal{D}$ none of which is equal to $\mathbb{I}$ so that $\mathbb{I}^{\prime}$ with representative ${\sf b}^{\mathbb{I}^{\prime}}:={\sf b}^{\mathbb{I}}+\sum_{i}\alpha_{i}{\sf b}^{\mathbb{I}_{i}}$, $0\not=\alpha_{i}\in\mathbb{F}$, is foldable.
Earlier, we defined $\kappa(\mathcal{D})$ as the number of $s$-complete interval modules in a decomposition $\mathcal{D}$ of a zigzag module $\mathbb{M}_{ZZ}$. Recall that some of the full interval modules in $\mathcal{D}$ may not be $s$-complete but can be converted to be $s$-complete. To account for those, we introduce the following notation.
Notation 6.1.
Let $\mathbb{M}_{ZZ}$ be an $s$-unfolding of $\mathbb{M}$ and $\mathcal{D}$ be any of its direct decomposition. Denote by $\tau(\mathcal{D})$ the number of interval modules $\mathbb{I}$ in $\mathcal{D}$ such that $\mathbb{I}$ is convertible and $\bar{\mathbb{I}}$ is invertible in $\mathcal{D}$. Observe that $\kappa(\mathcal{D})\leq\tau(\mathcal{D})$ because $s$-complete interval modules are convertible by condition (i) in Definition 6.1 and their complements are invertible by definition.
Proposition 6.2.
$\kappa(\mathcal{D})\leq{\sf{rk}}(\mathbb{M})\leq\tau(\mathcal{D})$.
Proof.
It follows from Proposition 4.1 that $\kappa(\mathcal{D})\leq{\sf{rk}}(\mathbb{M})$. So, we only show ${\sf{rk}}(\mathbb{M})\leq\tau(\mathcal{D})$. Observe that if ${\sf{rk}}(\mathbb{M})=0$ there is nothing to prove since $\tau(\mathcal{D})$ is non-negative by definition, so assume ${\sf{rk}}(\mathbb{M})\not=0$. Consider an $s$-complete decomposition $\mathcal{D}^{*}$ of $\mathbb{M}_{ZZ}$ which exists according to Proposition 4.1. Let $\mathbb{J}_{1},\ldots,\mathbb{J}_{r}$, $r={\sf{rk}}(\mathbb{M})$, denote the set of these $s$-complete modules in $\mathcal{D}^{*}$. We show by induction that there is a set of full modules $\mathbb{I}_{1},\ldots,\mathbb{I}_{r}$ in the direct decomposition $\mathcal{D}$ of $\mathbb{M}_{ZZ}$ so that $\mathbb{I}_{i}$ is convertible to $\mathbb{J}_{i}$ and $\bar{\mathbb{I}_{i}}$ invertible, $i\in[1,r]$ establishing the claim.
For the base case, consider $\mathbb{J}_{1}$. By Proposition 6.1, there is a set of limit modules $\mathbb{I}^{\prime}_{i}$ in $\mathcal{D}$ so that there exist unique $0\not=\alpha_{i}\in\mathbb{F}$ giving ${\sf b}^{\mathbb{J}_{1}}=\sum_{i}\alpha_{i}{\sf b}^{\mathbb{I}^{\prime}_{i}}$. Choose any full module among $\mathbb{I}_{i}^{\prime}$s as $\mathbb{I}_{1}$ which is guaranteed to exist by Proposition 6.1. Then, consider the decomposition $\mathcal{D}_{1}$ that only replaces $\mathbb{I}_{1}$ with $\mathbb{J}_{1}$ in $\mathcal{D}_{0}=\mathcal{D}$. Observe that $\mathbb{M}_{ZZ}=\mathbb{I}_{1}\oplus\overline{\mathbb{I}_{1}}=\mathbb{J}_{1}\oplus\overline{\mathbb{J}_{1}}=\mathbb{I}_{1}\oplus\overline{\mathbb{J}_{1}}$ establishing that $\overline{\mathbb{I}_{1}}\cong\overline{\mathbb{J}_{1}}$ where $\overline{\mathbb{J}_{1}}$ is invertible (actually foldable). This means $\overline{\mathbb{I}_{1}}$ is invertible as required.
To complete the induction, assume that, for $1\leq j<r$, we already have that $\mathbb{I}_{1},\ldots,\mathbb{I}_{j}$ are convertible to $\mathbb{J}_{1},\ldots,\mathbb{J}_{j}$ and $\overline{\mathbb{I}_{1}},\ldots,\overline{\mathbb{I}_{j}}$ are invertible. Also, we have the decompositions $\mathcal{D}_{0},\ldots,\mathcal{D}_{j}$ of $\mathbb{M}_{ZZ}$ where $\mathcal{D}_{i}$ is obtained inductively from $\mathcal{D}_{i-1}$ by replacing $\mathbb{I}_{i}$ with $\mathbb{J}_{i}$. By Proposition 6.1, we have ${\sf b}^{\mathbb{J}_{j+1}}=\sum_{i}\alpha_{i}{\sf b}^{\mathbb{I}^{\prime}_{i}}$, $\alpha_{i}\in\mathbb{F}$, for some limit modules in $\mathcal{D}_{j}$. We claim that the collection $\{\mathbb{I}^{\prime}_{i}\}$ includes a full module other than $\mathbb{J}_{1},\ldots,\mathbb{J}_{j}$. If not, we have ${\sf b}^{\mathbb{J}_{j+1}}=(\sum_{i}\gamma_{i}{\sf b}^{\mathbb{J}_{i}})+(\sum_{k}\gamma^{\prime}_{k}{\sf b}^{\mathbb{I}^{\prime}_{k}})$ for $\gamma_{i},\gamma^{\prime}_{k}\in\mathbb{F}$ where none of $\{\mathbb{I}^{\prime}_{k}\}$ is full. Then, ${\sf b}^{\mathbb{J}_{j+1}}-\sum_{i}\gamma_{i}{\sf b}^{\mathbb{J}_{i}}=\sum_{k}\gamma^{\prime}_{k}{\sf b}^{\mathbb{I}^{\prime}_{k}}$. The LHS gives a representative of an $s$-complete module which is a sum of representatives of limit modules(RHS) none of which is full contradicting Proposition 6.1. Thus, the set $\{\mathbb{I}^{\prime}_{i}\}$ includes a full module other than $\mathbb{J}_{1},\ldots,\mathbb{J}_{j}$. Let $\mathbb{I}_{j+1}$ in $\mathcal{D}_{j}$ be any such full module. Observe that $\mathbb{I}_{j+1}$ is also in $\mathcal{D}$. It follows that $\mathbb{I}_{j+1}$ in $\mathcal{D}$ is convertible to $\mathbb{J}_{j+1}$. Also, with the same reasoning as in the base case, one can check that $\overline{\mathbb{I}_{j+1}}$ in $\mathcal{D}$ is invertible. ∎
Theorem 6.1.
Let $\mathbb{M}_{ZZ}$ be an $s$-unfolding of $\mathbb{M}$ and $\mathcal{D}$ be any direct decomposition of $\mathbb{M}_{ZZ}$.
  1. 1.
    If every convertible full module $\mathbb{I}$ in $\mathcal{D}$ where $\overline{\mathbb{I}}$ is invertible is $s$-complete, then $\kappa(\mathcal{D})={\sf{rk}}(\mathbb{M})$ ($\mathcal{D}$ is $s$-complete) else
  2. 2.
    Let $\mathbb{I}$ be a convertible module where $\overline{\mathbb{I}}$ in $\mathcal{D}$ is invertible. Let $\mathcal{D}^{\prime}$ be the direct decomposition obtained from $\mathcal{D}$ by replacing $\mathbb{I}$ with the converted module $($$\mathrm{Convert}(\mathbb{I})$ in step 4 of GenRank)), then $\kappa(\mathcal{D}^{\prime})=\kappa(\mathcal{D})+1$.
Proof.
For (1), observe that, $\kappa(\mathcal{D})=\tau(\mathcal{D})$ in this case implying $\kappa(\mathcal{D})={\sf{rk}}(\mathbb{M})=\tau(\mathcal{D})$ due to Proposition 6.2.
For (2), observe that an interval module $\mathbb{J}\neq\mathbb{I}$ in $\mathcal{D}$ is foldable iff it is foldable in $\mathcal{D}^{\prime}$ because the only affected module in $\mathcal{D}$ is $\mathbb{I}$. We observe that $\overline{\mathbb{J}}$ remains invertible in $\mathcal{D}^{\prime}$ if and only if it were so in $\mathcal{D}$. It follows that, compared to $\mathcal{D}$, the decomposition $\mathcal{D}^{\prime}$ has exactly one more foldable module, namely $\mathrm{Convert}(\mathbb{I})$, with its complement being invertible. Hence $\kappa(\mathcal{D}^{\prime})=\kappa(\mathcal{D})+1$. ∎
The algorithm GenRank draws upon Theorem 6.1. To accommodate computations with a fixed finite precision, we assume that the field $\mathbb{F}$ is finite. The algorithm takes every full interval module $\mathbb{I}$ that is not foldable and checks if it is convertible. If it is convertible, it converts it and then check if its complement $\overline{\mathbb{I}}$ for invertibility. It continues converting such modules to foldable modules until it cannot find any to convert. The current direct decomposition $\mathcal{D}$ of $\mathbb{M}_{ZZ}$ changes with these conversions. At the end of this process, all convertible modules in the current decomposition whose complements are invertible become foldable themselves. Their number then coincides with ${\sf{rk}}(\mathbb{M})$. To determine the existence of the limit modules whose addition makes $\mathbb{I}$ foldable, we take the help of an annotation matrix $A_{p}$  [17, Chaper 4] for each complex $K_{p}$, $p\in P_{ZZ}$, computed in step 2. Without further elaborations, we only mention that annotation for a $k$-cycle $z\in Z_{k}(K_{p})$ (cycle group in degree $k$) is the coordinate of its class $[z]$ in a chosen basis of $H_{k}(K_{p})$; see [7] (annotation algorithm in [7] is described for $\mathbb{F}=\mathbb{Z}_{2}$ though it easily extends for any finite field $\mathbb{F}$). The representative cycles maintained for every interval module in the decomposition of $\mathbb{M}_{ZZ}$ at point $p$ form a cycle basis $z_{1}],\ldots,[z_{g}]$ of $H_{k}(K_{p})$. We will see later that testing a full module for foldability amounts to calculating the annotations of the representative cycles of limit modules for certain points $p\in P_{ZZ}$ and performing certain matrix reductions with them; see section 6.1.
Algorithm GenRank ($P$-filtration $F$, $k\geq 0$)
  • Step 1. Unfold $P$ and $F$ into a zigzag path $P_{ZZ}$ and a zigzag filtration $F_{ZZ}$ respectively
  • Step 2. Compute an annotation matrix $A_{p}$ for every complex $K_{p}$ in $F_{ZZ}$, $p\in P_{ZZ}$
  • Step 3. Compute a barcode for $F_{ZZ}$ with representative $k$-cycles; Let $\mathcal{I}$ denote the set of full interval modules corresponding to full bars in the current decomposition $\mathcal{D}$
  • Step 4. For every module $\mathbb{I}\in{\mathcal{I}}$ do
    • If $\mathbb{I}$ is convertible in $\mathcal{D}$, update $\mathcal{D}$ with $\mathbb{I}\leftarrow\mathrm{Convert}(\mathbb{I})$
    • If $\overline{\mathbb{I}}$ is invertible in $\mathcal{D}$, mark $\mathbb{I}$ complete
  • Output the number of complete interval modules/*may output converted modules*/
Now we focus on the crucial checks in step 4. Fix the degree of all homology groups to be $k\geq 0$, which form the vector spaces of the modules in our discussion.
6.1 Convertibility of $\mathbb{I}$ and computing $\mathrm{Convert}(\mathbb{I})$
Assume that $\mathbb{I}$ is a full interval module in the current direct decomposition $\mathcal{D}$ of $\mathbb{M}_{ZZ}=\bigoplus_{i}\mathbb{I}_{i}=\bigoplus_{i}\mathbb{I}^{[p_{b_{i}},p_{d_{i}}]}$. For each interval module $\mathbb{I}_{i}=\mathbb{I}^{[p_{b_{i}},p_{d_{i}}]}$ the algorithm computes a sequence of representative $k$-cycles $\{z^{i}_{p_{j}}\in Z^{k}(K_{p_{j}})\mid p_{j}\in\{p_{b_{i}},p_{b_{i}+1},\ldots,p_{d_{i}}\}\subseteq P_{ZZ}\}$. A representative for a limit module $\mathbb{I}^{[p_{b_{i}},p_{d_{i}}]}$ is given by the homology classes $z^{i}_{p_{b_{i}}}],\ldots,[z^{i}_{p_{d_{i}}}]$. Such a representative for each limit module could be computed by an adaptation of the zigzag persistence algorithm in [30] or in [17, Chapter 4,section 4.3]. Instead, we apply the more efficient algorithm given in [15]. Then, a conversion of a convertible module $\mathbb{I}$ updates this representative as we update $\mathbb{I}$ to $\mathrm{Convert}(\mathbb{I})$.
For any point $p\in P_{ZZ}$, call the set of all points that fold to $s(p)$ its partners and denote it as $\mathrm{prt}(p)$. We check if there are limit modules whose representatives when added to the representative of $\mathbb{I}$ makes the vectors at each point in $\mathrm{prt}(p)$ the same for every point $p\in P_{ZZ}$ because that converts $\mathbb{I}$ to a foldable interval module. The partner sets partition $P_{ZZ}$. We designate an arbitrary point, say $p$, in each partner set $\mathrm{prt}(p)$ where $|\mathrm{prt}(p)|>1$ as the leader of $\mathrm{prt}(p)$. Let $L$ denote the set of these leaders. For every point $p\in L$, we do the following.
Let $\mathbb{I}_{1},\ldots,\mathbb{I}_{\ell}$ denote the set of limit modules in the current decomposition $\mathcal{D}$ of $\mathbb{M}_{ZZ}$ and WLOG assume $\mathbb{I}=\mathbb{I}_{1}$. The module $\mathbb{I}$ is convertible iff for all $p\in L$, there exist $\alpha_{i}\in\mathbb{F}$ so that ${\sf b}^{\mathbb{I}_{1}}_{p}+\sum_{i=2}^{\ell}\alpha_{i}{\sf b}^{\mathbb{I}_{i}}_{p}=({\sf b}^{\mathbb{I}_{1}}_{p_{j}}+\sum_{i=2}^{\ell}\alpha_{i}{\sf b}^{\mathbb{I}_{i}}_{p_{j}})$ for every $p_{j}\in\mathrm{prt}(p)$. Written differently, for all $p\in L$ and for every $p_{j}\in\mathrm{prt}(p)$, we require
${\sf b}^{\mathbb{I}_{1}}_{p}+{\sf b}^{\mathbb{I}_{1}}_{p_{j}}=\sum_{i=2}^{\ell}\alpha_{i}({\sf b}^{\mathbb{I}_{i}}_{p}+{\sf b}^{\mathbb{I}_{i}}_{p_{j}})=\sum_{i=2}^{\ell}\alpha_{i}v^{i}_{j}.$ (8)
where $v^{i}_{j}:={\sf b}^{\mathbb{I}_{i}}_{p}+{\sf b}^{\mathbb{I}_{i}}_{p_{j}}$.
Figure 6: Illustration of matrix reduction with linearization to check Eq. (8); (top) $p$ is representative of $\mathrm{prt}(p)=\{p\}\cup\{p_{1},\ldots,p_{t}\}$, each row $\mathbb{I}_{i}$ is a limit module with annotations of representatives at points in $\mathrm{prt}(p)$; (bottom left) row $\sf I_{i}$ consists of annotations of cycles $\{\hat{z}^{i}_{p_{j}}\}_{j}$ where $\hat{z}^{i}_{p_{j}}=z^{i}_{p}+z^{i}_{p_{j}}$; (bottom middle) annotations of cycles in row $\sf I_{i}$ form matrix $A_{i}^{p}$ in the stacks of matrices; (bottom right) each matrix $A_{i}^{p}$, $i=2,\ldots,\ell$, linearized into a vector and combined into a larger matrix $A^{p}$ and $A_{1}^{p}$ into vector $\sf v_{p}$.
Our goal is to determine coefficients $\alpha_{i}$s if Eq. (8) holds. To do this, for every $i=1,\ldots,\ell$ and every $p\in L$, consider the matrices $A_{i}^{p}$ whose columns are annotations of the representative cycles $\hat{z}_{p_{j}}^{i}$ of $v^{i}_{j}$ for every $p_{j}\in\mathrm{prt}(p)$; see Figure 6 for an illustration. We can compute the $\mathbb{F}$-sum of two chains $\hat{z}_{p_{j}}^{i}=z_{p_{j}}^{i}+z_{p}^{i}$ and add their annotation vectors which is well defined because their annotation vectors have the same dimension $g_{p}=\dim H_{k}(K_{p})$ since all complexes at $p$ and $p_{j}$s were made equal during unfolding. Notice that, if $t_{p}=|\mathrm{prt}(p)|$, the matrix $A_{i}^{p}$ has dimensions $g_{p}\times t_{p}$. Then, checking Eq. (8) boils down to determining $\alpha_{i}$s so that $A_{1}^{p}=\sum_{i=2}^{\ell}\alpha_{i}A_{i}^{p}$. This can be done by using the linearization trick presented in [18]. Each of the $g_{p}\times t_{p}$ matrices $A_{i}^{p}$, $i\in\{2,\ldots,\ell\}$, is linearized into a vector $\sf v_{i}$ of length $t_{p}g_{p}$ and concatenated into a matrix of dimensions $t_{p}g_{p}\times\ell$. Let $A^{p}$ denote this matrix (Figure 6,bottom-right).
Suppose that there are $t=\sum_{p\in L}t_{p}$ points in $\cup_{p}\mathrm{prt}(p)$ and $g=\max_{p}\{g_{p}\}$. We need to check Eq. (8) simultaneously for all $p\in L$ to determine the $\alpha_{i}$s. To do this, we concatenate matrices $A^{p}$ for all $p\in L$ each of dimensions $t_{p}g_{p}\times\ell$ to create a matrix $A$ of dimensions $(\sum_{p}g_{p}t_{p})\times\ell$. Then, a vector $\sf v$ of dimension $\sum_{p}g_{p}t_{p}$ is created by concatenating vectors $\sf v_{p}$ of dimension $t_{p}g_{p}$ obtained by linearizing the matrices $A_{1}^{p}$ for all $p\in L$. Checking if Eq. (8) holds simultaneously for all $p\in L$ boils down to checking if $\sf v$ is in the column space of $A$. This is a matrix rank computation on a matrix of dimensions $(\sum_{p}g_{p}t_{p})\times\ell$ which can be done in $O(\ell^{\omega-1}\sum_{p}g_{p}t_{p})=O(\ell^{\omega-1}gt)$ time where $\omega<2.373$ is the matrix multiplication exponent [22].
If Eq. (8) holds, we need to compute $\alpha=[\alpha_{2},\ldots,\alpha_{\ell}]^{T}$ which can be done by solving the linear system $A\alpha={\sf v}$. Then, we update $\mathbb{I}$ to $\mathrm{Convert}(\mathbb{I})$ with the representative ${\sf b}^{\mathbb{I}}+\sum_{i=2}^{\ell}\alpha_{i}{\sf b}^{\mathbb{I}_{i}}$ to get the new decomposition $\mathcal{D}$. This takes $O(\ell^{\omega-1}gt)$ time again.
Notice that $g=O(n)=O(t)$. There are $O(t)$ interval modules in the computed decomposition of $\mathbb{M}_{ZZ}$ and there are at most $O(n)$ full interval modules among them. So, we check at most $O(n)$ full interval modules for convertibility. If there are $\ell=O(t)$ number of limit modules, each convertibility check takes $O(\ell^{\omega-1}tg)=O(t^{\omega}n)$ time giving a total of $O(t^{\omega}n^{2})$ time for converting all convertible interval modules in step 4.
To finish the analysis of step 4, we need to consider the time spent for creating the annotations for cycle representatives of the limit modules. For a representative cycle $z_{p}$, we get its annotation in the complex $K_{p}$ by adding the annotations of length $g_{p}=\dim H_{k}(K_{p})$ for every simplex in $z_{p}$. Thus, to compute the annotations of all $g_{p}$ representative cycles at point $p$ (classes of non-trivial representative cycles at $p$ form a basis of $H_{k}(K_{p})$), we consider a $g_{p}\times n_{p}$ matrix $E$ whose columns are annotations for each of the $n_{p}$ $k$-simplices in $K_{p}$ and another matrix $G$ of dimensions $n_{p}\times g_{p}$ whose columns represent each cycle representative in the chain space of $K_{p}$. The product matrix $EG$ provides the annotations for each of the cycle representatives. This takes $O(n_{p}g_{p}^{\omega-1})=O(ng^{\omega-1})$ time. Accouting for all $p\in P_{ZZ}$ and $g=O(n)$, we obtain a complexity of $O(tn^{\omega})$. Then, over all iterations of convertibility, we get a time $O(tn^{\omega+1})$.
6.2 Invertibility of $\overline{\mathbb{I}}$
To check invertibility of $\overline{\mathbb{I}}$, we do the following: first, we observe that $\mathrm{Fld}_{s}(\mathbb{I})$ is well defined because $\mathbb{I}$ is already converted to be foldable. Consider the quotient module $\mathbb{Q}=\mathbb{M}/(\mathrm{Fld}_{s}(\mathbb{I}))$. If $\mathbb{Q}$ is a submodule of $\mathbb{M}$, we know that $\mathrm{Fld}_{s}(\mathbb{I})$ is a summand. In other words, $\overline{\mathbb{I}}$ is then invertible due to Theorem 3.1(3). We construct a $P$-filtration $\hat{F}$ that induces $\mathbb{Q}$ and check algorithmically if it is a submodule of $\mathbb{M}$ as follows.
After converting $\mathbb{I}$ to a foldable module, suppose that we obtain the representative cycles $\{[z_{p}]\}_{p\in P_{ZZ}}$ for $\mathbb{I}$. We cone these cycles in the $P$-filtration $F$ of the simplicial complex $K$ that induces $\mathbb{M}$. It means that, we use a dummy vertex $\omega$ and for every simplicial complex $K_{\mathrm{Fld}_{s}(p)}$, $p\in P_{ZZ}$, we introduce the simplex $\sigma\cup{\omega}$ for every $\sigma\in z_{p}$ if it is not already present. Let the new $P$-filtration be $\hat{F}$ which filters the coned complex denoted $\hat{K}$. The $P$-module induced by $\hat{F}$ is the quotient module $\mathbb{Q}$. We unfold $\mathbb{Q}$ implicitly by unfolding $\hat{F}$ to $\hat{F}_{ZZ}$ onto $P_{ZZ}$. Let $\mathbb{Q}_{ZZ}$ denote the zigzag module obtained by this unfolding process. We compute a direct decomposition $\mathcal{D}$ of $\mathbb{Q}_{ZZ}$ from $\hat{F}_{ZZ}$ by the zigzag persistence algorithm in [15]. We say $\mathcal{D}$ is extendable if there is a direct decomposition $\mathcal{D}^{\prime}$ of $\mathbb{M}_{ZZ}$ so that $\mathcal{D}^{\prime}=\mathcal{D}\cup\{\mathbb{I}\}$.
Proposition 6.3.
$\overline{\mathbb{I}}$ is invertible if and only if $\mathcal{D}$ is extendable.
Proof.
If $\mathcal{D}$ is extendable, we have $\mathbb{M}_{ZZ}=\mathbb{I}\oplus\mathbb{Q}_{ZZ}$ by definition of extendability and the fact that $\mathcal{D}$ is a direct decomposition of $\mathbb{Q}_{ZZ}$. It follows that $\overline{\mathbb{I}}$ is invertible becasue $\mathbb{Q}_{ZZ}$ is foldable by construction.
For the opposite direction, assume that $\overline{\mathbb{I}}$ is invertible. Since $\mathbb{I}$ is foldable, $\mathrm{Fld}_{s}(\mathbb{I})$ exists and is a summand of $\mathbb{M}$ according to Theorem 3.1(3). The quotient $\mathbb{Q}=\mathbb{M}/(\mathrm{Fld}_{s}(\mathbb{I}))$ is a submodule of $\mathbb{M}$. This means the morphisms $\mathbb{Q}(p\leq_{P}q)$ is a restriction of $\mathbb{M}(p\leq_{P}q)$ for every pairs $p\leq_{P}q$. Therefore, by the definition of unfolding, every morphism $\mathbb{Q}_{ZZ}(p\leq_{P_{ZZ}}q)$ is a restriction of $\mathbb{M}_{ZZ}(p\leq_{P_{ZZ}}q)$. We claim that $\mathcal{D}^{\prime}=\mathcal{D}\cup{\mathbb{I}}$ is a direct decomposition of $\mathbb{M}_{ZZ}$ establishing that $\mathcal{D}$ is extendable.
To prove the claim, observe that, due to the quotienting, the vector generating $\mathbb{I}(p)$ cannot lie in the vector space $\mathbb{Q}_{ZZ}(p)$ for every $p\in P_{ZZ}$. Then, $\mathbb{M}_{ZZ}(p)=\mathbb{I}(p)\oplus\mathbb{Q}_{ZZ}(p)$ for every $p\in P_{ZZ}$. Also, for every summand $\mathbb{I}^{\prime}$ of $\mathbb{Q}_{ZZ}$ in $\mathcal{D}$, the morphism $\mathbb{I}^{\prime}(p\leq_{P_{ZZ}}q)$ is a restriction of $\mathbb{M}_{ZZ}(p\leq_{P_{ZZ}}q)$ because we already established that $\mathbb{Q}_{ZZ}(p\leq_{P_{ZZ}}q)$ is a restriction of $\mathbb{M}_{ZZ}(p\leq_{P_{ZZ}}q)$. Therefore, for every interval module $\mathbb{I}_{i}\in\mathcal{D}\cup\{\mathbb{I}\}$, every morphism $\mathbb{I}_{i}(p\leq_{P_{ZZ}}q)$ is a restriction of $\mathbb{M}_{ZZ}(p\leq_{P_{ZZ}}q)$ over all pairs $p\leq_{P_{Z}Z}q$ while $\mathbb{M}_{ZZ}(p)=\oplus_{i}\mathbb{I}_{i}(p)$. It follows that $\mathcal{D}^{\prime}$ is a direct decomposition of $\mathbb{M}_{ZZ}$. ∎
We test if $\mathcal{D}$ is extendable by checking if the linear maps connecting two vectors (cycle classes) $z_{p}]$ and $z_{q}]$ for all immediate pairs $p,q\in P_{ZZ}$ in every interval module in $\mathcal{D}$ are preserved in $\mathcal{D}^{\prime}$. This means that if $p\leq_{P_{ZZ}}q$, the cycles $z_{p}$ and $z_{q}$ which are homologous in the complex $\hat{K}_{q}$ remain homologous in the complex $K_{q}$. This can be easily determined by checking the annotations of the cycles $z_{p}$ and $z_{q}$ in $K_{q}$. We need to compute the annotations of representative cycles for the interval modules in $\mathcal{D}$ in the complex $K_{q}$. We have already seen that it costs $O(tn^{\omega})$ time in total. Thus, the extendability of $\mathcal{D}$ can be checked in $O(tn^{\omega})$ time.
Time complexity. Let the graph $G=(P,E(p))$ of the poset $P$ indexing the input filtration $F$ in GenRank have a total of $m$ vertices and edges. Recall that $K_{p}$ denote the complex at point $p\in P$. Let $e_{p,q}=|K_{p}\setminus K_{q}|$ where $(p,q)$ is a directed edge in $G$, that is, $e_{p,q}$ is the number of simplices that are added in the filtration going from $p$ to $q$. Let
$e=\sum_{(p,q)\in E(P)}e_{p,q}\mbox{ and }t=\max\{m,e\}.$ (9)
The quantity $t$ is an upper bound on the size of the input $P$-filtration and the size of the poset $P$. Let every complex $K_{p}$ for every point $p\in P$ has at most $n$ simplices. Then, step 1 of GenRank takes at most $O(t)$ time to unfold the poset to $P_{ZZ}$ with the procedure described in section 5.1. Producing the zigzag filtration $F_{ZZ}$ takes at most $O(t)$ time. Step 2 takes at most $O(n^{3})$ time to produce the annotation matrix [17] for every complex $K_{p}$, $p\in P_{ZZ}$, giving a total of $O(tn^{3})$ time. Step 3 takes $O(t^{2}n)$ time to compute the zigzag barcode with representatives by the algorithm in [15].
Step 4 for testing convertibility and converting of $\mathbb{I}$ takes a total of $O(t^{\omega}n^{2})$ time as discussed before. Testing invertibility of $\bar{\mathbb{I}}$ for every $\mathbb{I}$ involves coning out the filtration $F$ and unfolding $\hat{F}$ which cannot take more than $O(t)$ time. Computing the zigzag barcode of $\hat{F}_{ZZ}$ with representative cycles takes $O(t^{2}n)$ time by the algorithm in [15] giving a total time of $O(t^{2}n^{2})$ for $O(n)$ full interval modules like $\mathbb{I}$. The verification of the extendability of the interval decomposition $\mathcal{D}$ of the quotient module $\mathbb{Q}_{ZZ}$ takes $O(tn^{\omega})$ time, giving a total time of $O(tn^{\omega+1})$ over all. Therefore, step 4 takes a total time $O(t^{2}n^{2}+tn^{\omega+1})$. Accounting for all terms, we get:
Theorem 6.2.
Let $F$ be an input $P$-filtration of a complex with $n$ simplices where $F$ and $P$ have size at most $t$. The algorithm GenRank computes ${\sf{rk}}(\mathbb{M})$ in $O(tn^{\omega+1}+t^{\omega}n^{2})$ time where $\mathbb{M}$ is the $P$-module induced by $F$. With $n=O(t)$, the bound becomes $O(t^{\omega+2})$.
Proof.
The correctness of GenRank follows from Theorem 6.1 because step 4 effectively either increases the count $\kappa(\mathcal{D})$ or determines that $\kappa(\mathcal{D})$ cannot be increased in which case $\mathcal{D}$ is $s$-complete. The time complexity claim follows from our analysis. ∎
7 Special case of degree-$d$ homology for $d$-complexes
In this section, we show that when a $P$-module $\mathbb{M}$ is induced by applying the homology functor in degree $d\geq 0$ on a $P$-filtration of a $d$-dimensional simplicial complex, we have a much more efficient algorithm for computing ${\sf{rk}}(\mathbb{M})$. The key observation is that, in this case, the representatives for the interval modules in a direct decomposition of $\mathbb{M}_{ZZ}$ takes a special form. In particular, the following Proposition holds leading to Theorem 7.1.
Proposition 7.1.
Let $\mathbb{M}$ be a $P$-module where $\mathbb{M}(p)=H_{d}(K_{p})$ with $K_{p}$ being a simplicial $d$-complex. For an interval module $\mathbb{I}^{[p_{b},p_{d}]}$ in a direct decomposition of $\mathbb{M}_{ZZ}$ and for any two points $p_{i},p_{j}\in[p_{b},p_{d}]$, the representative $d$-cycles $z_{p_{i}}$ and $z_{p_{j}}$ are the same.
Proof.
First assume that $p_{i}$ and $p_{j}$ are immediate points in $P_{ZZ}$ and without loss of generality let $p_{i}\rightarrow p_{j}$. According to the definition of representatives, we must have the homology classes $z_{p_{i}}]$ and $z_{p_{j}}]$ homologous in $K_{p_{j}}$. Since $K_{p_{j}}$ is a $d$-complex, $z_{p_{i}}]=[z_{p_{j}}]$ only if $z_{p_{i}}=z_{p_{j}}$ as chains. It follows by transitivity that this is true even if $p_{i}$ and $p_{j}$ are not immediate. ∎
It follows from the above proposition that a full interval module $\mathbb{I}$ in a direct decomposition of $\mathbb{M}_{ZZ}$ is foldable. Next proposition says that even the complement $\overline{\mathbb{I}}$ is invertible. Then, applying Theorem 3.1(3), we can claim that $\mathrm{Fld}_{s}(\mathbb{I})$ is a full interval summand of $\mathbb{M}$.
Proposition 7.2.
Let $\mathbb{I}$ be any full interval module in a direct decomposition of $\mathbb{M}_{ZZ}$ where $\mathbb{M}_{ZZ}$ is constructed as in Proposition 7.1. Then, $\mathbb{I}$ is foldable and $\overline{\mathbb{I}}$ is invertible.
Proof.
$\mathbb{I}$ is foldable due to Proposition 7.1. We need to show that $\overline{\mathbb{I}}$ is invertible.
For the full module $\mathbb{I}$, there is a single $d$-cycle that forms a fixed representative at each point in $P_{ZZ}$ (Proposition 7.1). Let $z$ be such a $d$-cycle and $\sigma$ be any $d$-simplex in $z$. Delete $\sigma$ from the complex $K_{p}$ for every $p\in P_{ZZ}$. Let $\mathbb{M}_{ZZ}^{-}$ denote the $P_{ZZ}$-module induced by the homology functor in degree $d$ on the zigzag filtration $F_{ZZ}^{-}$ obtained from the original filtration $F_{ZZ}$ by deleting the simplex $\sigma$ everywhere. It is easy to verify that $\mathbb{M}_{ZZ}^{-}$ is $s$-foldable because $K_{p}\setminus\sigma=K_{p^{\prime}}\setminus\sigma$ for every $p,p^{\prime}\in P_{ZZ}$ with $s(p)=s(p^{\prime})$. A direct decomposition of $\mathbb{M}_{ZZ}$ is obtained from a direct decomposition of $\mathbb{M}_{ZZ}^{-}$ by adding a full bar with the representative $z$ or equivalently the summand $\mathbb{I}$. It follows that $\overline{\mathbb{I}}$ is invertible. ∎
Theorem 7.1.
Let $\mathbb{M}$ be a module constructed as in Proposition 7.1 from a $P$-filtration $F$ of a $d$-complex where $P$ and $F$ have size at most $t$. Then, ${\sf{rk}}(\mathbb{M})$ is the number of full interval modules in any direct decomposition of $\mathbb{M}_{ZZ}$ which can be computed in $O(t^{\omega})$ time.
Proof.
By Proposition 7.2 and Theorem 3.1(3), we get that $\kappa(\mathcal{D})=\tau(\mathcal{D})$ for any direct decomposition $\mathcal{D}$ of $\mathbb{M}_{ZZ}$. Then, it follows from Proposition 6.2 that ${\sf{rk}}(\mathbb{M})=\kappa(\mathcal{D})$, which is exactly equal to the number of full interval modules in any direct decomposition of $\mathbb{M}_{ZZ}$. Therefore, ${\sf{rk}}(\mathbb{M})$ can be simply obtained by computing the zigzag barcode of $\mathbb{M}_{ZZ}$ (no need of computing the representatives). This can be done in $O(t^{\omega})$ time with the fast zigzag algorithm [14]. ∎
If $\mathbb{M}$ is induced by a $P$-filtration of a graph, then ${\sf{rk}}(\mathbb{M})$ can be computed even faster. Observe that every $1$-cycle that represents a full bar must continue to be present from the initial graph $G_{p_{0}}$ at $p_{0}\in P$ to the final graph $G_{p_{m}}$ at $p_{m}\in P$. This suggests the following algorithm. Take $G_{p_{m}}$; delete all edges and vertices that are deleted as one moves along $P_{ZZ}$ from $p_{0}$ to $p_{m}$, but do not insert any of the added edges and vertices. In the final graph $G_{p_{m}}^{\prime}$ thus obtained (which may be different from $G_{p_{m}}$ because we ignore the inserted edges and vertices along $P_{ZZ}$), we compute the number of independent $1$-cycles. This number can be computed by a depth first search in $G_{p_{m}}^{\prime}$ in linear time. Then, as an immediate corollary we have Theorem 7.2.
Theorem 7.2.
If $\mathbb{M}$ is induced by degree-$1$ homology of a $P$-filtration $F$ of a graph with a total of $n$ vertices and edges, then ${\sf{rk}}(\mathbb{M})$ can be computed in $O(n+t)$ time where $P$ and $F$ have size at most $t$.
8 Conclusions and discussions
Analyzing a mutliparameter persistence module with the help of one parameter persistence modules is not new. It has been introduced in the context of computing matching distances [2, 10, 23] for $2$-parameter modules. The unfolding/folding technique proposed here offers a different slicing technique. By producing one zigzag path instead of multiple slices, the unfolding preserves the structural maps of the original module $\mathbb{M}$ in a lossless manner. We showed how to use them for reconstructing full interval modules and hence for computing the generalized rank. It will be interesting to see what other invariants one can reconstruct using the folding/unfolding of persistenec modules. A natural candidate would be to compute the limit and colimit of the original module from its zigzag straightening. Recent advances in zigzag persistence computations [14, 30, 32] can then be taken advantage of for computing limits and colimits.
A natural question to ask is that if the approach laid out in this paper works even when the original input module is unfolded into another module that is not necessarily a zigzag module. This could lead to a more generalized theory for unfolding. Unfortunately, this question does not have a positive answer. The main bottleneck arises while attempting to extend Proposition 6.1. This proposition says that any full interval module that is a submodule of a zigzag module $\mathbb{M}_{ZZ}$ is a sum of the limit modules in a direct decomposiiton of $\mathbb{M}_{ZZ}$. We cannot extend this result to every module $\mathbb{M}$ that is not a zigzag module. The proof of Theorem 6.1 uses Proposition 6.2 which depends on this result in a fundamental way. To see the obstacle, consider the module $\mathbb{M}$ in Figure 3 (left,top). This module has a full interval module $\mathbb{I}$ as a submodule, namely $\mathbb{I}$ is given by the vector spaces spanned by the vector $v$ at every point and the internal map as the identity. However, the module $\mathbb{M}$ itself is not decomposable and hence has no limit module as a summand. Thus, the full module $\mathbb{I}$ has no expression as a sum of limit modules that are summands of $\mathbb{M}$. The situation does not improve even if $\mathbb{I}$ is a summand instead of being merely a submodule. The module $\mathbb{M}$ in Figure 3 (left,bottom) has a full interval module $\mathbb{I}$ as a summand whose vector spaces are spanned by the vector $u+v$ at every point. Also, a different decomposition of $\mathbb{M}$ can have a full interval module $\mathbb{I}^{\prime}$ as a summand whose vector spaces are spanned by the vector $u$ at every point. But, $\mathbb{I}$ cannot be expressed as a sum of limit modules in this decomposition which has $\mathbb{I}^{\prime}$ as the only limit module.
The special structure of the zigzag modules allows us to express any full submodule as a sum of limit modules that are summands, which in turn lets our algorithm to check if a full module in the current decomposition is convertible or not.
Acknowledgment.
We thank Iason Papadopoulos of University of Bremen for useful comments on the paper.